A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

28 December 2023 Version 1.13

Title word cross-reference

$17.27 [Whi23]. $48.99 [Whi23]. $97.00 [Whi23].

Abundance [ACAL21, CWLJ21, HDKC20, HDJ^{+}22, HRSD21, HSJ^{+}20, JSI^{+}21, MHO^{+}22, OGT22, GBR^{+}23, HJ23, VMS^{+}23]. Acid [RLO20]. Acknowledgments [Ano20g, Ano21g, Ano23g, RFR^{+}22c]. Acoustic [BEH^{+}22, BGR^{+}23, LBC^{+}23]. across [JZGE20, PMG20, PLRR21]. Actions [MH20]. Active [HGW^{+}21]. acute [LDWS23]. Adaptive [HOL21]. admixture [DEBW23]. Adult [BHF21, CWLJ21, DRR^{+}21, IZMP22, RFR^{+}22a, RFR^{+}22b]. Adulthood [FBW^{+}21, SHS^{+}21]. Advancing [CWKY22]. Affect [RAZ^{+}21]. Affected [GSN^{+}22, WBAB21]. Affects [NH21]. after [ADR^{+}23, BDW21, IRE^{+}22, WCSJ20]. Age [GSN^{+}22, MHO^{+}22, RLMCC21, THH^{+}22, RKAF23]. Age-0 [THH^{+}22].

Age-1 [MHO^{+}22, RKAF23]. Age-2 [GSN^{+}22]. Alabama [KDW^{+}23]. Alaska [DB23, HBH^{+}20, MWS^{+}23, OWB^{+}22, RMS^{+}22, RBS20]. albula [MVK23]. Allee [MVK23]. Alligator [FRF^{+}22]. Allometric [WRRG22]. along [AFB^{+}22, BSP^{+}22]. Alosa [LJR^{+}21]. Alteration [TM22]. Alterations [AMZC^{+}20]. Amargosa [PS20]. Amazonas [LCFSS21]. American [GBR^{+}23, GLSM^{+}21, LJ20, WSS^{+}22]. among [EECA20, FBW^{+}21, LJR^{+}21, PZTC20, RVY20, SHS^{+}21]. Anadromous [BHJ^{+}22]. Anadromy [QPS^{+}21]. Analgesic [RAZ^{+}21]. Analysis [BDW21, FRF^{+}22, SMC^{+}22, VCV23, WBKK23]. anchoita [CBF^{+}20]. Anchovy [CBF^{+}20]. Angler [MSH^{+}20, SMW20, MH23, VCV23]. anglers [VCV23]. Angling [DKS^{+}22, MSH^{+}20, MH23]. Annual [DT21]. Anomalously [TDB20]. Apache [DGG^{+}23]. Apalachicola [DKD^{+}22]. Appalachian [KLK^{+}22, JZH^{+}23, NH21]. Application [LJ20]. Applications [Bra20]. Approach [WRRG22]. Apps [BAF22]. Aquatic [MLZ^{+}20]. archival [EME23]. Arctic [TKM^{+}22]. Area [KSW21, MLZ^{+}20]. Areas [THH^{+}22]. Argentine [CBF^{+}20]. Arizona [BKM23]. artedi [EKO^{+}21, MOE^{+}23]. Artificial [BPLL21, KGP22, NMG^{+}21]. Aspects [JPC^{+}20]. Assemblage [CHDF20, PARM23]. Assemblages [KGP22, PSBGDRH20]. Assess [CWKY22, GBR^{+}23]. Assessing [HSJ^{+}20]. Assessment [Bra20, GBR^{+}21, KGP22, MHI^{+}20, SOK^{+}22]. Assign [CWKY22, MHI^{+}20]. Assignment [IBT^{+}20, WBKK23]. assist [RCC^{+}23]. Associated [BHC^{+}20, KSW21, KTJ^{+}23, YYK^{+}23]. Association [DERF20, HCS23]. Associations [FBW^{+}21, MLF21, SHS^{+}21, VBQ23]. At-Risk [FBW^{+}21, SHS^{+}21]. Atlantic [AFB^{+}22, CBF^{+}20, BHF21, BSP^{+}22, DCF^{+}23, FDDB21, HWK20, HDKC20, LT21, MHO^{+}22, MMPZ21, PWM^{+}21, RKAF23]. Attraction [TKM^{+}22]. Australian [SMM^{+}22]. Availability [DRR^{+}21, HLC^{+}23]. Avoidance [WJCA20].

Bad [LCM^{+}21]. barriers [GBR^{+}23]. Basal [PSBGDRH20]. Based [AMZC^{+}20, SN23]. Basin [CBB^{+}23, HBH^{+}20, LFL^{+}20, TKM^{+}22, EME23]. Bass [BSC^{+}21, DKS^{+}22, ERH^{+}20, HJ23, LCSNB23, MBD^{+}20, MLF21, PLRR21, PZTC20, RLO20, SMW20, VACS^{+}22, WBKK23]. Behavior [MLZ^{+}20, RFR^{+}22a, RFR^{+}22b, LBC^{+}23]. Behavioral [AMZC^{+}20, LT21]. Behaviors [SMW20]. benefits [RH23]. Benthic [BRJ^{+}22]. between [BHTH21, PWM^{+}21, VBQ23]. Big [TKM^{+}22]. bigheaded [TWP23, VMS^{+}23]. biochronology [DB23]. Bioeconomic [SMM^{+}22]. Bioenergetics [BDW21, JBO23, SN23]. bioenergetics-based [SN23]. Biological [Hax23, LBC^{+}23, RKAF23]. Biology [Tel20, VHA^{+}23]. Biomarkers [RLO20]. Black [MHO^{+}22, MLF21]. Blackwater [CHDF20]. Blooms [BHC^{+}20]. Blue [NF20, OGD22, RLMCC21, SHO21, BE23, OGD23]. Bluegill [MLZ^{+}20]. Body [JDZ^{+}21, Rey21]. Bonneville [BTM^{+}20]. Book [Whi23]. Boost [HDJ^{+}22]. bottlenecks [ADR^{+}23]. Boundaries [ERH^{+}20]. Brandon [SOK^{+}22]. Brazil [LCFSS21]. Brazos [TM22]. Broader [PWM^{+}21]. Broadscale [EMG^{+}22]. broodstock [PGT23]. Brook [BKVH20, EMG^{+}22, KTJ^{+}23, KLK^{+}22, KSW21, NKL^{+}20, NH21, SW22, VBQ23]. Brown [KSW21, OGT22]. Build [KGP22]. Bull [VBQ23]. Burbot [ARH^{+}21, MHI^{+}20]. burgeoning [DYG^{+}23]. Busycotypus [AFB^{+}22].

California [TDB20, SMC^{+}22]. Can [PARM23, RH23]. Canada [BT23, PWM^{+}21]. canal [GBR^{+}23, BVB20]. canaliculatus [AFB^{+}22]. Canyon [DYG^{+}23]. Capability [CWKY22]. Capacity [BHJ^{+}22, RCC^{+}23]. Capelin [RBS20]. captive [PGT23]. Captivity [JPC^{+}20]. Capture [SMW20]. Captured [MB20, MLZ^{+}20]. Carp [BEH^{+}22, WJCA20, TWP23]. carps [VMS^{+}23]. Case [LCM^{+}21, Wip21]. Catch [BAF22, LJC^{+}22, MSH^{+}20, BHC^{+}23]. Catch-and-Release [LJC^{+}22, MSH^{+}20, BHC^{+}23]. Catches [EECA20]. Catfish [JBD23, NF20, SHO21]. Catostomus [CBB^{+}23]. Cause [GLSM^{+}21, MVK23]. central [MWS^{+}23]. Champlain [IZMP22]. Change [ARH^{+}21, BDW21, GLSM^{+}21, KK23, MWS^{+}23]. Changes [LT21, WRRG22]. changing [YYK^{+}23]. Channel [JBD23]. Channeled [AFB^{+}22]. Channels [PPH^{+}21]. characteristics [Hax23, HCS23, MWS^{+}23]. Characterization [BVB20]. Characterizes [HRS^{+}20]. Chemistry [AKB^{+}20]. Chiapas [PSBGDRH20]. Chinook [BT23, EME23, FJ20, FDS^{+}20, GSN^{+}22, HLC^{+}23, JTB^{+}21, JTB^{+}22, JSI^{+}21, KSG^{+}22, LBC^{+}23, PO20, PGT23, PM23, PPH^{+}21, RMS^{+}22, RLO20, WBD^{+}22, WBAB21]. Chub [DERF20, DYG^{+}23, KK23, HCS23]. Chulitna [HBH^{+}20]. Chum [ACAL21]. Cichla [LCFSS21]. Cisco [EJ20, JZGE20, EKO^{+}21]. Ciscoes [EKO^{+}21, RVY20, MOE^{+}23]. Clark [HBH^{+}20]. clarkii [PS20]. Clearnose [NJM22]. Clearwater [CHDF20]. Climate [ACAL21, ARH^{+}21, BW22, GLSM^{+}21, JTB^{+}21, JTB^{+}22, DGG^{+}23, KK23, MWS^{+}23]. Climatic [MHO^{+}22]. Clip [FRF^{+}22]. Coast [BSP^{+}22, MB20, AFB^{+}22]. Coastal [CWLJ21, KSH^{+}22, LJR^{+}21, PLRR21]. Cochineal [MHN^{+}22]. Cod [PWM^{+}21]. Coexistence [IJF21]. Coho [BVB20, JSI^{+}21, LFL^{+}20]. cold [RH23]. cold-pool [RH23]. collapse [MVK23, YYK^{+}23]. Colonization [FDDB21]. Colorado [BKM23, LBAW21]. Columbia [FDS^{+}20, PO20, PM23, WBD^{+}22]. Come [KGP22]. Comment [FBW^{+}21, SHS^{+}21]. Commercially [MB20]. Common [AFRHSB^{+}22, BEH^{+}22, SM21, WJCA20]. communities [LLB23]. Community [BRJ^{+}22, BKM23]. Comparative [MLZ^{+}20]. Comparing [EECA20]. Comparison [HRSD21, MOE^{+}23, MHM21, RLMCC21, RCC^{+}23]. Comparisons [WRRG22]. Competitor [ACAL21]. Complete [HBH^{+}20]. Complex [PMG20, LCSNB23, MOE^{+}23, SMC^{+}22]. Composition [TDB20]. compositions [PM23]. Concentration [MHN^{+}22]. Condition [GYL^{+}21, JPC^{+}20, JSI^{+}21, MBD^{+}20, MHI^{+}20]. Conditions [SMW20, YYK^{+}23]. Conductivity [Rey21]. Confluence [RZGL20]. Connectivity [BTM^{+}20, LLB23]. Conservation [SSS^{+}22, PARM23].

Conspecifics [JTB^{+}21, JTB^{+}22]. constraints [SN23]. Constructed [HRS^{+}20]. Consumption [LN21, MBD^{+}20, SHO21]. Contacts [EJ20]. Contemporary [BHTH21]. Contrasts [CHDF20]. Contributions [CWLJ21]. Copepod [ACAL21]. Coregonus [EKO^{+}21, MVK23, MOE^{+}23]. Counts [HRSD21]. Crab [OGD22, OGD23]. crabs [BE23]. Crappie [BSC^{+}21]. Crayfish [PS20]. Cren [GYL^{+}21]. Croaker [LT21, MHO^{+}22]. Cryptic [HSJ^{+}20]. Cue [HMJW20]. cues [DCF^{+}23]. Current [GBR^{+}21, TDB20]. Cutthroat [BTM^{+}20, BHTH21, DEBW23, HLC21, LBAW21, MH20, OK22]. Cyanobacteria [BHC^{+}20]. Cycle [RHA21].

Dace [BW22, SMC^{+}22]. Daily [LJR^{+}21, RHA21]. Dam [FBW^{+}21, HDW^{+}22, HPM^{+}21, SOK^{+}22, SHS^{+}21, KDW^{+}23]. Dams [IRE^{+}22]. Darter [HSJ^{+}20]. Data [HOL21, LJ20]. Day [TR20]. decline [BKM23]. Declines [CWLJ21, YYK^{+}23, KTJ^{+}23]. Decomposition [KTG^{+}20]. Decrease [LCFSS21]. Dee [DCF^{+}23]. Defining [IBT^{+}20]. Delays [MMPZ21]. Delta [AFRHSB^{+}22, PPH^{+}21]. Demographics [IBT^{+}20, WCSJ20]. Denil [TKM^{+}22]. Densities [DRR^{+}21]. Density [CBF^{+}20, HLC21, NMG^{+}21, NF20, PS20]. Density-Dependence [NF20]. Density-Dependent [HLC21, PS20]. Density-Independent [HLC21]. Dependence [NF20]. Dependent [HLC21, PS20]. Depends [BKVH20]. Depletion [LJ20]. Depth [BPLL21]. depths [EME23]. Derived [NKL^{+}20]. Describing [IBT^{+}20]. Detectably [GSN^{+}22]. Detected [CFL^{+}21]. Detection [GBR^{+}21, KSW21, KTJ^{+}23]. Determined [CWKY22]. Deterring [BEH^{+}22]. Development [BSP^{+}22, JHM^{+}21]. Diadromous [Wip21]. Did [ALH21, RAZ^{+}21]. Diet [TDB20]. dietary [HLC^{+}23]. Diets [BRJ^{+}22, NH21, TWP23]. differ [DCF^{+}23]. differences [PARM23]. Differential [PRB^{+}20]. Differentially [LFL^{+}20]. Discharge [LFL^{+}20]. Dispersal [HRS^{+}20]. Distance [ERH^{+}20]. Distinct [IJF21, SMC^{+}22]. Distribution [JTB^{+}21, JTB^{+}22, KSW21, KSH^{+}22, PGT23, RBS20, GBR^{+}23]. Distributions [FDS^{+}20]. Disturbed [HBH^{+}20]. Dive [HRSD21]. Divergence [CFL^{+}21, OWB^{+}22]. Divergent [CFL^{+}21]. Diverse [Nie21]. Diversity [BT23, OWB^{+}22, PSBGDRH20, JBD23, LLB23]. DNA [GBR^{+}23, KTJ^{+}23, KGP22]. Do [SM21]. Documenting [Ken22]. Dogfish [MB20]. Don’t [Rai22]. Downstream [HMJW20]. Drainage [ERH^{+}20, DB23]. Drainages [ERH^{+}20, RKAF23]. drier [DGG^{+}23]. Driftless [KSW21]. driven [BE23]. Drivers [HWK20, LJR^{+}21]. Drum [LDWS23, MHO^{+}22, TWP23]. Drying [PLRR21]. dual [DCF^{+}23]. dual-spawning [DCF^{+}23]. duration [LBC^{+}23]. during [EME23, FDDB21, MVK23, RAZ^{+}21, RZGL20, YYK^{+}23]. Dwelling [OK22]. Dye [MHN^{+}22]. Dynamics [DERF20, FJ20, IBT^{+}20, LBAW21, OGT22, THH^{+}22, MH23, SN23].

Early [TDB20, VACS^{+}22]. Eastward [GBR^{+}21]. Eat [ALH21]. ebook [Whi23]. Ecology [AFRHSB^{+}22, IJF21, SPVM20, FBK^{+}23, Whi23]. ectoparasites [KTJ^{+}23]. Eel [LJ20, GBR^{+}23]. Eels [WSS^{+}22]. effect [HLC^{+}23, LDWS23, MVK23]. Effectively [BSP^{+}22]. Effects [ARH^{+}21, AKB^{+}20, BRJ^{+}22, BSC^{+}21, DKS^{+}22, DKD^{+}22, HLC21, JPC^{+}20, JHM^{+}21, KSG^{+}22, LCFSS21, MBD^{+}20, MSH^{+}20, NMG^{+}21, NF20, PS20, Rai22, TWP23, DEBW23, JZH^{+}23, LCSNB23, RKAF23]. Efficiency [TKM^{+}22]. Effort [BAF22, MH23]. Electro [RFR^{+}22a, RFR^{+}22b]. Electro-Immobilization [RFR^{+}22a, RFR^{+}22b]. Electroanesthesia [RAZ^{+}21]. Electrofishing [HRSD21]. electronically [WZS^{+}23]. elevated [BHC^{+}23]. Elvers [LJ20]. Elwha [QPS^{+}21]. Embryonic [JHM^{+}21]. Embryos [ARH^{+}21, KTG^{+}20, MHN^{+}22, PKZW23]. Emphasis [EJ20]. Empirical [LJ20]. encountered [EME23]. Endangered [BHC^{+}20, HWK20, PARM23]. Endemic [EKO^{+}21, HCS23]. Energetic [LT21]. Energy [CBF^{+}20, OK22]. Engraulis [CBF^{+}20]. Entrance [TKM^{+}22]. Environmental [Bra20, CHDF20, KGP22, LJR^{+}21, RMS^{+}22, SMW20, GBR^{+}23, KTJ^{+}23, LBC^{+}23, RKAF23]. Equations [JDZ^{+}21]. Erie [CEL^{+}20, FBK^{+}23]. Erratum [Ano22a, JTB^{+}22, RFR^{+}22a]. error [WZS^{+}23]. Essington [Whi23]. estimate [MH23]. Estimates [CEL^{+}20, LCM^{+}21, SHO21]. Estimating [JBD23, RLMCC21, Rey21, RCC^{+}23]. Estimation [HRSD21]. Estuaries [OGT22]. Estuarine [LT21, SN23]. Estuary [MHO^{+}22, BE23, LBC^{+}23, SPVM20, WBD^{+}22]. Eulachon [BGR^{+}23]. Evaluate [MH20]. Evaluating [BHC^{+}23, DEBW23, HDW^{+}22, LBC^{+}23, LCM^{+}21]. Evaluation [HMJW20, NKL^{+}20, SSM20, JZH^{+}23, RCC^{+}23]. Events [RZGL20]. Evidence [OGD22, WCSJ20]. Example [HSJ^{+}20, HBH^{+}20, IBT^{+}20]. Exert [SM21]. Expansion [GBR^{+}21]. Experience [MSH^{+}20]. experiment [HLC^{+}23]. Experimental [PS20]. Explained [PRB^{+}20]. Exploitation [OK22, OGD23]. Exploring [AMZC^{+}20, BHTH21, MHO^{+}22]. Exposure [VACS^{+}22, LDWS23]. Expressions [BTM^{+}20]. Extreme [SW22].

face [MWS^{+}23]. Factors [DT21, HLC21, KSW21, HJ23, LBC^{+}23, RKAF23]. Fall [FJ20]. Family [HRS^{+}20]. Fast [WBD^{+}22]. Fatty [RLO20]. features [HCS23]. Federally [DERF20]. Feeding [WBD^{+}22, WBAB21]. Feeling [DRR^{+}21]. Female [BPLL21, JPC^{+}20, BE23]. Fin [FRF^{+}22, SOK^{+}22]. Fine [PWM^{+}21]. Fine-Scale [PWM^{+}21]. Finescale [BW22]. Fire [JTB^{+}21, JTB^{+}22]. First [Wal20]. Fish [ALH21, AKB^{+}20, BDW21, BKVH20, BEH^{+}22, CHDF20, FBW^{+}21, HOL21, HDW^{+}22, HRSD21, KGP22, LCFSS21, MHM21, PSBGDRH20, RAZ^{+}21, Rey21, SOK^{+}22, SHS^{+}21, BKM23, KK23, KDW^{+}23, LLB23, PARM23, SN23]. Fisheries [HBH^{+}20, Nie21]. Fishery [OGT22, SMM^{+}22]. Fishes [HSJ^{+}20, Tay20, Tel20, Wip21]. Fishing [BAF22, Ken22, LJC^{+}22, LCFSS21, SMM^{+}22, SMW20, WZS^{+}23, SMM^{+}22].

Fishways [TKM^{+}22]. fjord [LBC^{+}23]. flooded [KK23]. Floodplain [PPH^{+}21]. Florida [BSP^{+}22, DKD^{+}22, HJ23, PLRR21]. Flow [Bra20, PWM^{+}21]. Flows [SW22]. Fluctuation [ARH^{+}21]. Following [HPM^{+}21, RFR^{+}22a, RFR^{+}22b]. Food [SHO21, HLC^{+}23, SN23]. Foodscape [OK22]. Foraging [DCB^{+}21, MHM21, SN23]. Form [EKO^{+}21]. forums [VCV23]. Four [RLMCC21]. Fox [THH^{+}22]. Fraser [BT23, BGR^{+}23]. Frequencies [HOL21, JDZ^{+}21]. Frequency [HWK20]. Freshwater [LDWS23, OWB^{+}22, TWP23, BT23]. future [DGG^{+}23, KK23, KK23].

Gametogenesis [MHI^{+}20]. Gaps [DCB^{+}21]. Gar [FRF^{+}22]. Gene [PWM^{+}21]. Generalized [LJ20]. Genetic [AFB^{+}22, CEL^{+}20, CFL^{+}21, ERH^{+}20, HRS^{+}20, IBT^{+}20, NKL^{+}20, OWB^{+}22, SSM20, CBB^{+}23, JBD23, ALH21]. Genetics [KLK^{+}22, WCSJ20, KDW^{+}23]. Genomic [BVB20, JZH^{+}23, SMC^{+}22]. Geometric [WRRG22, MOE^{+}23]. Geomorphology [JTB^{+}21, JTB^{+}22]. Ghats [AKB^{+}20]. gill [KTJ^{+}23]. GIS [NKL^{+}20]. GIS-Derived [NKL^{+}20]. Gizzard [JDZ^{+}21, VMS^{+}23]. Goby [BRJ^{+}22, GBR^{+}21]. Gradient [LT21]. Gradients [JZGE20]. Grand [DYG^{+}23]. Grande [ADR^{+}23]. gravel [LLB23]. Grayling [TKM^{+}22]. Great [CBB^{+}23, DCF^{+}23]. Greenland [GYL^{+}21]. Groundwater [RZGL20]. Growth [ACAL21, BDW21, BRJ^{+}22, DT21, GLSM^{+}21, HLC21, JSI^{+}21, MBD^{+}20, NF20, Rai22, WBD^{+}22, DB23, MWS^{+}23]. Guiding [HMJW20]. Guild [AKB^{+}20]. Gulf [KGP22, DCB^{+}21, DKD^{+}22, PWM^{+}21]. Gulkana [RMS^{+}22]. Guts [ALH21].

Habitat [BHF21, BKVH20, BSC^{+}21, DCB^{+}21, DRR^{+}21, DERF20, ERH^{+}20, FDDB21, HSJ^{+}20, HRS^{+}20, MLF21, PMG20, PLRR21, PRB^{+}20, RHA21, HJ23, HCS23, JBD23, RCC^{+}23, SN23]. Habitat-Linked [PRB^{+}20]. Habitats [NF20, WBD^{+}22, DGG^{+}23]. Halibut [GYL^{+}21, HDKC20]. happy [VCV23]. Hard [RLMCC21, KDW^{+}23]. hard-part [KDW^{+}23]. hardcover [Whi23]. Harlequin [HSJ^{+}20]. Harvest [NH21]. Harvested [MB20]. Harvester [HDKC20]. Hatchery [ADR^{+}23, EMG^{+}22, GSN^{+}22, KSH^{+}22, TR20, HLC^{+}23, PGT23, PM23]. Hatchery-Origin [KSH^{+}22, PM23]. Hatchery-Raised [GSN^{+}22]. Head [HDW^{+}22]. Herring [DRR^{+}21, LJR^{+}21]. High [BHJ^{+}22, CFL^{+}21, RZGL20, THH^{+}22]. High-Occupancy [THH^{+}22]. High-Resolution [BHJ^{+}22]. High-Temperature [RZGL20]. Histology [CWKY22]. Historic [CWLJ21]. History [BTM^{+}20, FBW^{+}21, OWB^{+}22, RBS20, SHS^{+}21, VACS^{+}22, BT23]. Hole [TKM^{+}22]. Holt [Ken22]. Hood [BVB20]. Hook [WJCA20]. Hudson [BHF21]. Humpback [DB23, DYG^{+}23]. Huron [EKO^{+}21, MOE^{+}23]. Hurricane [DKD^{+}22]. Hybrid [MBD^{+}20]. Hybridization [EJ20, FDS^{+}20, LCSNB23]. Hybrids [EKO^{+}21, WRRG22]. Hydroacoustic [JDZ^{+}21]. Hydrodynamics [RZGL20]. Hydrologic [TM22]. Hydropower [ARH^{+}21, WSS^{+}22]. hyperoxia [LDWS23]. Hypomesus [MHN^{+}22]. Hypoxia [BSC^{+}21, MBD^{+}20, LDWS23]. hypoxia/hyperoxia [LDWS23].

Ice [LJC^{+}22]. Idaho [VBQ23]. Identification [ALH21, THH^{+}22, KTJ^{+}23]. identifies [SMC^{+}22]. Identifying [BW22, DCB^{+}21]. If [KGP22]. Illinois [SOK^{+}22]. Imaging [HOL21]. Immobilization [RFR^{+}22a, RFR^{+}22b, Rey21]. Impact [TM22]. Impacts [AMZC^{+}20, TDB20, KK23, KDW^{+}23]. Implantation [RFR^{+}22a, RFR^{+}22b]. Implications [ARH^{+}21, AKB^{+}20, BTM^{+}20, EECA20, OGT22, SSS^{+}22, WBKK23]. Importance [BPLL21, PSBGDRH20, RHA21, VACS^{+}22]. Including [EKO^{+}21, RLMCC21]. Inclusive [Nie21]. Incorporating [HDKC20]. Increases [CWLJ21, ADR^{+}23, LLB23]. Incubation [WBAB21]. Independent [HLC21]. Index [HDKC20, SN23, GYL^{+}21]. India [AKB^{+}20]. Indices [ACAL21, RMS^{+}22]. Individual [AMZC^{+}20, BEH^{+}22]. Individual-Based [AMZC^{+}20]. Induced [AMZC^{+}20, WJCA20]. Induces [KTG^{+}20, PKZW23]. infer [MH23]. Inferred [RLO20]. Influence [HJ23, JTB^{+}21, JTB^{+}22, LFL^{+}20, SW22, SMW20, VMS^{+}23, JBD23, LBC^{+}23, MWS^{+}23]. Influences [TR20]. Influencing [DT21]. inform [PARM23]. Information [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, DCB^{+}21, LN21]. inland [Hax23]. Insights [BHJ^{+}22]. Instantaneous [GLSM^{+}21]. Intake [OK22]. Integrated [MH20]. Intensities [MHM21]. Intensity [NH21]. Interacting [HLC21]. Interactive [MBD^{+}20]. Interior [FJ20]. Intermittent [BTM^{+}20]. International [Wal20]. Intervals [BHF21]. Intracoelomic [RFR^{+}22a, RFR^{+}22b]. Introduced [KSW21, VBQ23]. Introduction [BDW21, Whi23]. Introgression [BKVH20, EMG^{+}22]. Invasive [BEH^{+}22, HGW^{+}21, MHM21, PS20, SHO21, PKZW23, VMS^{+}23]. Inventory [HBH^{+}20]. Invertebrate [BRJ^{+}22]. investigate [WZS^{+}23]. Investigation [AKB^{+}20]. Investment [PZTC20]. Iowa [KSW21]. Isotope [FRF^{+}22, RLO20]. Issue [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f]. Issues [LJC^{+}22].

J [Ken22]. Japanese [MHN^{+}22]. Joaquin [PPH^{+}21]. John [TR20]. Juvenile [BSP^{+}22, BHC^{+}20, DCB^{+}21, DRR^{+}21, GYL^{+}21, HMJW20, IZMP22, JSI^{+}21, LT21, MBD^{+}20, NF20, PPH^{+}21, RLO20, SM21, WBD^{+}22, DB23, MWS^{+}23, RCC^{+}23, VBQ23].

Kenai [MWS^{+}23]. Kennebec [Wip21]. Key [CBF^{+}20]. keystone [KK23]. Kids [ALH21]. Killifish [SM21]. kinematics [LDWS23]. Kiyi [VHA^{+}23]. Klamath [BHC^{+}20]. Know [Rai22]. Knowledge [HDKC20].

Lahontan [BHTH21]. Lake [BRJ^{+}22, BHC^{+}20, FBK^{+}23, Hax23, HGW^{+}21, HBH^{+}20, HRS^{+}20, IBT^{+}20, IRE^{+}22, IJF21, IZMP22, JZGE20, KTG^{+}20, MSH^{+}20, PKZW23, SPVM20, THH^{+}22, WCSJ20, AMZC^{+}20, CEL^{+}20, EKO^{+}21, ERH^{+}20, FBK^{+}23, HJ23, HGW^{+}21, IBT^{+}20, IZMP22, JHM^{+}21, KTG^{+}20, MOE^{+}23, RVY20, VHA^{+}23]. Lakes [DRR^{+}21, LLB23]. Lamprey [CWLJ21, HMJW20, HPM^{+}21, HDJ^{+}22]. Lampreys [ALH21]. Landscape [MWS^{+}23]. Lane [WBD^{+}22]. Large [DT21, IJF21, MHO^{+}22, BE23, PARM23]. Large-Scale [MHO^{+}22]. Largemouth [BSC^{+}21, DKS^{+}22, HJ23, LCSNB23, PLRR21, PZTC20, SMW20]. Larval [ALH21, AMZC^{+}20, HRS^{+}20, TWP23]. Learning [WJCA20]. Length [FBW^{+}21, OGT22, SHS^{+}21, RKAF23]. leuciscid [HCS23]. level [HCS23]. Levels [CFL^{+}21, MB20]. lice [KTJ^{+}23]. Lidocaine [RAZ^{+}21]. Life [BTM^{+}20, HDJ^{+}22, OWB^{+}22, RBS20, VACS^{+}22, WBD^{+}22, BT23]. Lifelong [Ken22]. Light [HMJW20, MHM21, RHA21]. Lighting [NMG^{+}21]. Limit [DRR^{+}21]. Limitation [OGD22, SN23]. Lineage [LCSNB23]. lineages [SMC^{+}22]. Linked [ACAL21, JSI^{+}21, PRB^{+}20]. lipid [HLC^{+}23]. Little [DERF20]. Lived [WCSJ20]. Local [GYL^{+}21, RAZ^{+}21]. Lock [HDW^{+}22, SOK^{+}22, KDW^{+}23]. lock-and-dam [KDW^{+}23]. Long [BKM23, DERF20, MLZ^{+}20, WCSJ20]. Long-Lived [WCSJ20]. Long-Standing [MLZ^{+}20]. Long-Term [DERF20, BKM23]. Lost [BHC^{+}20]. Louis [SPVM20]. Louisiana [OGD22, OGD23]. Low [HDW^{+}22, KDW^{+}23]. Low-Head [HDW^{+}22]. Low-Use [HDW^{+}22, KDW^{+}23]. lower [BGR^{+}23, THH^{+}22].

Main [WBD^{+}22]. Main-Stem [WBD^{+}22]. Maine [MMPZ21, PWM^{+}21, RKAF23]. Maintaining [BTM^{+}20]. Make [Nie21]. makes [VCV23]. male [HLC^{+}23]. Males [MVK23]. Management [MH20, NKL^{+}20]. Marine [LCM^{+}21, TDB20]. Marking [MHN^{+}22]. Maryland [LN21]. Massachusetts [LJR^{+}21, MB20]. matches [WZS^{+}23]. matching [WZS^{+}23]. Maternal [JHM^{+}21]. Maturation [GSN^{+}22, WBAB21, HLC^{+}23]. mature [BE23, DB23]. May [GLSM^{+}21, MVK23]. Measuring [HOL21]. Mediates [NF20]. Menominee [IRE^{+}22]. Mesocosms [BHC^{+}20]. Metapopulation [BTM^{+}20]. Methods [Bra20, BSP^{+}22, CWKY22, GBR^{+}21, RFR^{+}22a, RFR^{+}22b, RCC^{+}23]. Methylmercury [AMZC^{+}20, MB20]. Methylmercury-Induced [AMZC^{+}20]. Mexican [AFRHSB^{+}22]. Mexico [KGP22, PSBGDRH20]. Michael [DKD^{+}22]. Michigan [AMZC^{+}20, ERH^{+}20, IBT^{+}20, IRE^{+}22]. Microchemistry [SOK^{+}22, KDW^{+}23]. microhabitat [HCS23]. Mid [MHO^{+}22]. Mid-Atlantic [MHO^{+}22]. Middle [LCFSS21]. Midwestern [EMG^{+}22]. Migrant [HMJW20]. Migration [LFL^{+}20, RFR^{+}22a, RFR^{+}22b, BE23, BGR^{+}23, DCF^{+}23, EME23, LBC^{+}23]. Migrations [LJR^{+}21]. Migratory [IJF21, DEBW23, WBKK23]. miniature [LBC^{+}23]. minijack [PGT23]. Minijacks [GSN^{+}22]. Minimally [HBH^{+}20]. Minnow [ADR^{+}23]. Minnows [TM22]. Mississippi [ERH^{+}20]. mixed [WBKK23]. mixed-stock [WBKK23]. Mobile [BAF22]. Model [AMZC^{+}20, FRF^{+}22, LJ20, MH20, JBO23]. Modeling [RMS^{+}22]. models [MH23]. Molecular [CBB^{+}23, KTJ^{+}23]. Mollusk [SMM^{+}22]. Montana [TKM^{+}22]. Montane [RZGL20]. Morphological [JZGE20]. Morphology [PRB^{+}20, TDB20]. Morphometric [WRRG22]. morphometrics [MOE^{+}23]. Mortality [BHC^{+}20, GLSM^{+}21, KTG^{+}20, BHC^{+}23, PKZW23]. Mountains [KLK^{+}22]. Movement [BEH^{+}22, FJ20, FDDB21, IRE^{+}22, IJF21, LBAW21, MMPZ21, PMG20, THH^{+}22]. Movements [IZMP22, RZGL20]. Multifaceted [BTM^{+}20]. Multipiscivore [MHM21]. Multiple [JDZ^{+}21]. Multiscale [AKB^{+}20]. Muskellunge [BHC^{+}23, SPVM20]. mykiss [LCM^{+}21].

Natal [FJ20]. National [HBH^{+}20]. Native [GLSM^{+}21, KSW21, MHM21, RVY20, SOK^{+}22, BKM23, JZH^{+}23, KK23]. Natural [GLSM^{+}21, KSH^{+}22, PO20, WCSJ20, PM23]. Natural- [KSH^{+}22]. Natural-Origin [PO20]. Naturally [FJ20, KSH^{+}22]. Nearshore [OGT22]. Need [IBT^{+}20]. Needs [LJC^{+}22]. Negro [LCFSS21]. Neighboring [FJ20]. Nevada [BHJ^{+}22, BKM23]. Newly [HRS^{+}20]. Niche [RVY20]. Night [NMG^{+}21]. Nipissing [JHM^{+}21]. nipponensis [MHN^{+}22]. Nisqually [LCM^{+}21]. Nonlethal [FRF^{+}22, MHI^{+}20]. Nonnative [BDW21, NF20, RVY20, RLO20, SM21, VACS^{+}22, DEBW23, JBO23]. Nonprobability [BAF22]. Nonregulated [AKB^{+}20]. Northern [LN21, KGP22, TDB20]. Northwest [FJ20, HDKC20, KSH^{+}22]. Novel [FDDB21, WSS^{+}22].

Occupancy [THH^{+}22]. Ocean [CBF^{+}20]. off [MB20]. Ohio [PZTC20]. Okeechobee [HJ23]. Oncorhynchus [LCM^{+}21]. Oneida [BRJ^{+}22]. Ontario [Hax23, JHM^{+}21]. Ontogenetic [LT21]. Opposing [DB23]. Optimization [SMM^{+}22]. Oregon [CWLJ21, BHC^{+}20, LFL^{+}20]. Organic [KTG^{+}20]. Origin [KSH^{+}22, PO20, PM23, WBKK23]. Origins [QPS^{+}21]. other [HJ23]. Otolith [MHN^{+}22, DB23]. Otoliths [RLMCC21]. Output [SSS^{+}22]. Overlap [FDS^{+}20, IZMP22]. Oxford [Whi23].

Pacific [CWLJ21, FBW^{+}21, FJ20, HPM^{+}21, HDJ^{+}22, KSH^{+}22, SHS^{+}21]. Paddlefish [SSS^{+}22]. pages [Whi23]. paperback [Whi23]. Parameters [BPLL21]. Parentage [WBAB21]. Parents [ALH21]. Park [HBH^{+}20]. Parr [FJ20, KSH^{+}22, RKAF23]. part [KDW^{+}23]. Partitioning [AFRHSB^{+}22, RVY20]. Passage [FBW^{+}21, HDW^{+}22, SOK^{+}22, SHS^{+}21, TKM^{+}22, DEBW23, WSS^{+}22]. Patterns [FDDB21, IJF21, LJR^{+}21, LFL^{+}20, OGT22, SSS^{+}22, BKM23, BGR^{+}23]. Pavon [LCFSS21]. Pee [DCF^{+}23]. Pelagophilic [TM22]. Pellet [KTG^{+}20]. Perch [AMZC^{+}20, BRJ^{+}22, CFL^{+}21]. Performance [PRB^{+}20, LDWS23]. Phenology [VACS^{+}22]. phylogenetic [CBB^{+}23]. Physiological [DKS^{+}22, JPC^{+}20]. Pike [SM21]. Piscataquis [MMPZ21]. Piscivore [BDW21]. pit [LLB23]. Plaines [SOK^{+}22]. planning [RCC^{+}23]. Plasticity [EJ20]. Polyodon [SSS^{+}22]. Pomoxis [PRB^{+}20]. pool [RH23]. Population [AFB^{+}22, BW22, BTM^{+}20, BHTH21, CWKY22, DERF20, EMG^{+}22, HSJ^{+}20, KLK^{+}22, LCFSS21, MH20, NJM22, OGT22, PWM^{+}21, RHA21, SSM20, SMC^{+}22, ADR^{+}23, CBB^{+}23, DYG^{+}23, KDW^{+}23, MVK23, YYK^{+}23]. Populations [BVB20, BKVH20, HLC21, PS20, PZTC20, Hax23, PM23, RH23]. Postrelease [RFR^{+}22b, RFR^{+}22a]. Potamodromous [IJF21]. Potential [BDW21, TDB20, ADR^{+}23, JBD23, LCSNB23]. Precocious [GSN^{+}22, WBAB21, HLC^{+}23]. Predation [NMG^{+}21]. Predator [ALH21, BHTH21, IJF21, NMG^{+}21]. Predatory [SM21]. Predict [LN21]. Predicting [DCB^{+}21]. Predictors [BPLL21, CBF^{+}20]. Predominant [OWB^{+}22]. Preserve [HBH^{+}20]. Press [Whi23]. Pressure [SM21]. Previous [CEL^{+}20]. Prey [BDW21, BHTH21, MHM21, YYK^{+}23]. Private [WJCA20]. Procambarus [PS20]. Proceedings [Wal20]. processes [MH23]. Produced [FJ20, KSH^{+}22]. production [HLC^{+}23, YYK^{+}23]. production-scale [HLC^{+}23]. Productivity [JZGE20, LCM^{+}21]. Profitable [Ken22]. Program [HGW^{+}21, PGT23]. projected [MWS^{+}23]. Protected [MLZ^{+}20]. Published [LN21]. Puget [ACAL21, LCM^{+}21]. Pupfish [PS20]. Pursuit [Ken22]. Putative [EKO^{+}21, CBB^{+}23].

Quality [BSC^{+}21]. quantify [KDW^{+}23]. Quantifying [EECA20, OK22]. Quantitative [Whi23].

RAD [CEL^{+}20]. RAD-Seq [CEL^{+}20]. radio [EME23]. radio-archival-tagged [EME23]. Rainbow [CWKY22, DT21, EECA20, OWB^{+}22, PKZW23, RVY20, YYK^{+}23]. Raised [GSN^{+}22]. Range [GBR^{+}21, GLSM^{+}21]. Ranges [OK22]. Rangewide [NJM22]. rare [SN23]. Rates [HLC21, PO20, SHO21, DYG^{+}23, MWS^{+}23, OGD23]. Ratios [EECA20]. Ray [SOK^{+}22]. Razorback [PMG20]. Reach [VBQ23]. Reach-scale [VBQ23]. Really [LCM^{+}21]. Rearing [RLO20]. Recolonization [HPM^{+}21]. Reconditioned [JPC^{+}20]. Reconstruction [BTM^{+}20, HRS^{+}20]. Recovery [Wip21]. recreational [MH23]. Recruitment [LJ20, SW22]. Red [BSP^{+}22, BPLL21, PS20]. Redds [JTB^{+}21, JTB^{+}22]. Redside [JBO23]. reducing [HLC^{+}23]. Reefs [KGP22]. Refines [CEL^{+}20]. Reflect [OWB^{+}22]. Refuge [PLRR21]. refuges [RH23]. Regime [JTB^{+}21, JTB^{+}22, WBAB21]. Region [PWM^{+}21, JZH^{+}23]. regulate [MH23]. Regulated [AKB^{+}20, DT21]. Regulates [RZGL20]. Reintroduced [BHTH21]. Reintroduction [BHJ^{+}22]. Relates [ERH^{+}20]. Relation [MHO^{+}22]. Relations [RHA21]. Relationship [BHTH21]. Relationships [HSJ^{+}20, CBB^{+}23]. Relative [CWLJ21, PSBGDRH20, PPH^{+}21]. Release [LJC^{+}22, MSH^{+}20, BHC^{+}23]. released [PGT23]. Relict [BW22]. Remnant [BTM^{+}20]. Removals [HPM^{+}21]. Repeat [EECA20, JPC^{+}20]. Replacement [EKO^{+}21]. reported [WZS^{+}23]. Reproduction [HGW^{+}21, WCSJ20, KK23]. Reproductive [BPLL21, CWKY22, MHI^{+}20, PZTC20, SSS^{+}22, Tel20, VHA^{+}23]. Research [LJC^{+}22]. Reservoir [BSC^{+}21, MLF21, PMG20, PZTC20, BKM23, PARM23, YYK^{+}23]. residence [LBC^{+}23]. Resident [OWB^{+}22, MWS^{+}23, VBQ23]. Resiliency [DGG^{+}23]. Resolution [BHJ^{+}22, HOL21]. Resource [AFRHSB^{+}22, OK22]. Resources [PSBGDRH20]. Response [DKS^{+}22, FBW^{+}21, MSH^{+}20]. Responses [BEH^{+}22, LT21, MH20]. Restoration [AKB^{+}20, RCC^{+}23]. Results [FDS^{+}20]. Resumption [QPS^{+}21]. Revealed [GYL^{+}21, BGR^{+}23, DB23]. Reveals [OK22]. Review [BAF22, Whi23]. Reviewer [Ano20g, Ano21g, Ano23g, RFR^{+}22c]. Richness [ACAL21, AKB^{+}20]. Rio [ADR^{+}23]. Riparian [NH21]. Risk [FBW^{+}21, SHS^{+}21]. River [AFRHSB^{+}22, BT23, DRR^{+}21, DB23, DKD^{+}22, DT21, HBH^{+}20, IRE^{+}22, KK23, LLB23, LJR^{+}21, LFL^{+}20, LCFSS21, MWS^{+}23, MMPZ21, PSBGDRH20, PMG20, PPH^{+}21, RMS^{+}22, SPVM20, SOK^{+}22, TKM^{+}22, THH^{+}22, WBD^{+}22, WBKK23, GBR^{+}23, BHF21, BHC^{+}20, BKM23, BGR^{+}23, DCF^{+}23, DERF20, EME23, ERH^{+}20, FDS^{+}20, HWK20, HDJ^{+}22, KSG^{+}22, KDW^{+}23, LBAW21, LCM^{+}21, PO20, QPS^{+}21, TR20, TM22, VMS^{+}23, VACS^{+}22, Wip21]. River-of-origin [WBKK23]. Riverine [BHF21, PPH^{+}21]. Rivers [AKB^{+}20, BHJ^{+}22, CWLJ21, FJ20]. Road [SOK^{+}22]. Robust [HPM^{+}21]. Rotenone [PKZW23]. Round [BRJ^{+}22, GBR^{+}21]. route [BE23]. Run [DRR^{+}21, RMS^{+}22, WBAB21]. Runs [EECA20].

Sacramento [CFL^{+}21, PPH^{+}21]. Safe [WSS^{+}22]. Salinity [LT21, MHO^{+}22]. Salish [Tay20]. Salmon [ACAL21, BVB20, BT23, EME23, FBW^{+}21, FJ20, FDS^{+}20, FDDB21, GSN^{+}22, HLC^{+}23, JTB^{+}21, JTB^{+}22, JSI^{+}21, KSG^{+}22, LBC^{+}23, LFL^{+}20, MMPZ21, PO20, PPH^{+}21, QPS^{+}21, RMS^{+}22, RLO20, RKAF23, SHS^{+}21, WBD^{+}22, WBAB21, MWS^{+}23, PGT23, PM23, RCC^{+}23]. Salmonid [NMG^{+}21, Rai22, RZGL20, RH23]. Salmonids [BHJ^{+}22, RH23]. Sample [BSP^{+}22]. Sampling [BAF22]. San [PPH^{+}21]. Sandhills [HCS23]. Saugeye [BSC^{+}21]. save [RH23]. Scale [JSI^{+}21, MHO^{+}22, PWM^{+}21, HLC^{+}23, VBQ23]. Scaling [SSS^{+}22]. Sea [HMJW20, Tay20]. Season [EECA20]. Seasonal [DCF^{+}23, FBK^{+}23, GYL^{+}21, IZMP22, MSH^{+}20, OGT22, PSBGDRH20, SW22]. Secondary [EJ20]. segregation [JBD23]. Selected [Ken22]. selective [DEBW23]. sentiment [VCV23]. Seq [CEL^{+}20]. Severity [PLRR21]. Sex [CWKY22, EECA20, MHI^{+}20]. Shad [GLSM^{+}21, JDZ^{+}21, VMS^{+}23]. Shallow [EKO^{+}21]. Shallow-Water [EKO^{+}21]. Shape [WRRG22]. Shifts [GLSM^{+}21]. Shiner [JBO23]. Short [RAZ^{+}21]. Short-Term [RAZ^{+}21]. Shrimp [OGT22]. Sidney [Ken22]. Sierra [BHJ^{+}22]. Signals [MHO^{+}22]. Significant [BHTH21]. Silvery [ADR^{+}23]. Simulated [DKS^{+}22]. Simulation [RHA21, RH23]. Sire [GSN^{+}22]. Sites [BW22]. Size [DRR^{+}21, JDZ^{+}21, JSI^{+}21, KSH^{+}22, LCFSS21, LCSNB23]. Skate [NJM22]. Small [KSH^{+}22]. Smallmouth [ERH^{+}20, RLO20, VACS^{+}22]. Smart [SMM^{+}22]. Smelt [EECA20, MHN^{+}22, RVY20]. Smolt [TR20]. Smolts [MMPZ21]. Snake [HDJ^{+}22, TR20]. Snakehead [LN21, Wal20]. Snapper [BSP^{+}22, BPLL21]. Snook [AFRHSB^{+}22, SM21]. Social [WJCA20]. Sockeye [QPS^{+}21]. Sonar [HOL21]. Sound [ACAL21, LCM^{+}21]. Sources [CFL^{+}21]. south [MWS^{+}23]. south-central [MWS^{+}23]. Southern [KLK^{+}22]. Southwestern [HLC21, WRRG22, CBF^{+}20, OWB^{+}22, SPVM20]. sp [CBB^{+}23]. Space [HDW^{+}22, MH23]. spathula [SSS^{+}22]. Spatial [FDS^{+}20, IZMP22, JTB^{+}21, JTB^{+}22, OGD23, OK22, SPVM20, SN23, FBK^{+}23, PARM23]. Spawn [LFL^{+}20]. Spawner [JSI^{+}21]. Spawning [BVB20, BHF21, CWKY22, EECA20, FDS^{+}20, HWK20, HRS^{+}20, JPC^{+}20, KSH^{+}22, LJR^{+}21, OGD22, RBS20, ADR^{+}23, DCF^{+}23, EME23, OGD23, PM23]. Speciation [EJ20]. Species [ACAL21, ALH21, AKB^{+}20, CBF^{+}20, FBW^{+}21, HOL21, MLF21, PRB^{+}20, SHS^{+}21, WCSJ20, WRRG22, BKM23, CBB^{+}23, KK23, SMC^{+}22]. Speckled [LCFSS21, SMC^{+}22]. Sperm [OGD22]. Spiny [MB20]. Spot [MHO^{+}22]. Spotfin [DERF20]. Spotted [FRF^{+}22]. spp [LJR^{+}21]. Spring [FJ20, FDS^{+}20, GSN^{+}22, WBAB21, BT23]. Spring-Run [WBAB21]. Squeeze [DRR^{+}21]. St. [SPVM20]. Stable [FRF^{+}22, RLO20]. Stages [HDJ^{+}22]. Standing [MLZ^{+}20]. State [LCFSS21, MH23]. state-space [MH23]. States [VCV23]. Status [HSJ^{+}20, LJC^{+}22]. Steelhead [JPC^{+}20, KSH^{+}22, LCM^{+}21, PO20, TR20, TDB20, WBD^{+}22, PM23]. Stem [WBD^{+}22]. Stimuli [BEH^{+}22]. Stock [IBT^{+}20, OGD22, ADR^{+}23, OGD23, WBKK23]. Stocked [BKVH20, LBAW21]. Stocked-Fish [BKVH20]. Stocking [WCSJ20, JZH^{+}23]. Stories [Nie21]. Strategies [PMG20, SMM^{+}22]. Stray [PO20, PM23]. Straying [QPS^{+}21, TR20, PGT23]. Stream [HRSD21, KSH^{+}22, OK22, MWS^{+}23, VBQ23]. Stream-Dwelling [OK22]. stream-resident [MWS^{+}23, VBQ23]. Streams [CHDF20, LJR^{+}21]. Strength [JDZ^{+}21]. Strength-to-Body [JDZ^{+}21]. Stress [DKS^{+}22]. Striped [MBD^{+}20, WBKK23]. Stroboscopic [BEH^{+}22]. Strong [HSJ^{+}20]. Structure [AFB^{+}22, BPLL21, BTM^{+}20, CEL^{+}20, CWKY22, EMG^{+}22, ERH^{+}20, NJM22, PWM^{+}21, SSM20, BKM23, PARM23]. Structures [RLMCC21, KDW^{+}23]. Structuring [NKL^{+}20]. Study [LCM^{+}21, RHA21, Wip21]. Sturgeon [BHF21, DCB^{+}21, DCF^{+}23, DKD^{+}22, HWK20, HRS^{+}20, IRE^{+}22, IZMP22, MSH^{+}20, THH^{+}22, WCSJ20]. Subarctic [RBS20]. subyearling [LBC^{+}23]. Success [MHM21, SMW20, VACS^{+}22, WCSJ20]. Sucker [PMG20, WRRG22, CBB^{+}23, PARM23]. Suckers [BHC^{+}20, RLMCC21, TKM^{+}22]. Suggests [CFL^{+}21]. suitability [SN23]. Summer [FDS^{+}20, BHC^{+}23, BT23]. Superior [RVY20, SPVM20, VHA^{+}23]. Supplementation [KSG^{+}22, ADR^{+}23]. Support [LJ20]. Supporting [PSBGDRH20]. Suppression [HGW^{+}21]. Surprise [CBB^{+}23]. Survey [BSP^{+}22, EECA20, LJ20]. Surveys [BAF22]. Survival [FBW^{+}21, JHM^{+}21, LBAW21, LCM^{+}21, MMPZ21, PLRR21, PPH^{+}21, RFR^{+}22a, RFR^{+}22b, SHS^{+}21, TDB20]. Sustainable [Ken22]. Swamp [PS20]. Swimming [EME23, LDWS23]. Sympatric [PRB^{+}20]. Symposium [Wal20]. Synchrony [IJF21]. System [HWK20, MLF21, DKD^{+}22, GBR^{+}23].

tagged [EME23]. Tailbeat [HOL21]. Tailrace [HDW^{+}22]. tailwater [YYK^{+}23]. Tanana [DB23]. Target [JDZ^{+}21]. Tautog [SSM20]. telemetry [BGR^{+}23]. Teleost [RAZ^{+}21, Tel20]. Tell [Nie21]. temensis [LCFSS21]. Temperature [ARH^{+}21, DKS^{+}22, LFL^{+}20, MBD^{+}20, MHO^{+}22, NF20, Rai22, RZGL20, WBAB21, BHC^{+}23, LDWS23]. temperatures [EME23]. Temporal [FDS^{+}20, OK22, OGD23, PARM23]. Tennessee [DERF20, VMS^{+}23]. Term [DERF20, RAZ^{+}21, BKM23]. Texas [OGT22]. Their [GLSM^{+}21, WRRG22, EME23]. Thermal [RZGL20, VACS^{+}22, RH23]. Threatened [DERF20]. Threats [RLO20]. Three [HOL21, MLF21, RKAF23, SMC^{+}22]. Threshold [Rey21]. Throughout [GLSM^{+}21]. Timber [NH21]. Time [MHN^{+}22, PPH^{+}21]. Timing [BE23, BHF21, LFL^{+}20, RMS^{+}22]. Timothy [Whi23]. Tools [MHI^{+}20]. Top [IJF21]. Tournament [SMW20]. Tracking [WCSJ20]. traditional [MOE^{+}23]. Transborder [PWM^{+}21]. Transfer [IRE^{+}22]. Translocation [BW22, CFL^{+}21]. Translocations [HDJ^{+}22]. Transmitter [RFR^{+}22a, RFR^{+}22b]. transmitters [LBC^{+}23]. Transportation [TR20]. Trap [IRE^{+}22]. Travel [PPH^{+}21]. Treatment [MHN^{+}22]. Trends [CWLJ21, MHO^{+}22, DB23]. Tributary [RZGL20]. trips [WZS^{+}23]. Trophic [AFRHSB^{+}22, PSBGDRH20]. Trophy [LCFSS21]. Tropical [PSBGDRH20]. Trout [BKVH20, BTM^{+}20, BHTH21, CWKY22, DEBW23, DT21, EMG^{+}22, FBK^{+}23, HGW^{+}21, HLC21, KTJ^{+}23, KLK^{+}22, KSW21, KTG^{+}20, LBAW21, MH20, NKL^{+}20, NH21, OWB^{+}22, OK22, PKZW23, SW22, TKM^{+}22, VBQ23, YYK^{+}23, DGG^{+}23]. turbine [WSS^{+}22]. Twenty [WCSJ20]. Twenty-Two [WCSJ20]. Two [BHJ^{+}22, CFL^{+}21, FBW^{+}21, FJ20, IRE^{+}22, RFR^{+}22a, RFR^{+}22b, SHS^{+}21, TDB20, WCSJ20, WRRG22]. Typical [EKO^{+}21].

U.S. [AFB^{+}22]. UK [Whi23]. Umpqua [LFL^{+}20]. Underwater [HRSD21]. United [VCV23]. Units [NKL^{+}20]. University [Whi23]. upper [DB23, PM23, BHC^{+}20, ERH^{+}20, FDS^{+}20, KSG^{+}22, PO20]. Upwelling [RZGL20]. urbanization [KK23]. USA [CWLJ21, CBB^{+}23, MMPZ21, PWM^{+}21]. Use [BHF21, FDDB21, GBR^{+}23, HDW^{+}22, HSJ^{+}20, HRS^{+}20, PPH^{+}21, KDW^{+}23]. Using [AMZC^{+}20, BAF22, HSJ^{+}20, KDW^{+}23, LN21, MH20, MH23, RMS^{+}22, RFR^{+}22a, RFR^{+}22b, SOK^{+}22, WRRG22, DB23, IBT^{+}20, LBC^{+}23, MOE^{+}23]. Usumacinta [AFRHSB^{+}22].

Validating [WZS^{+}23]. Validation [FRF^{+}22]. Valley [CBB^{+}23]. Variable [MHM21]. Variation [BEH^{+}22, JZGE20, LJR^{+}21, PSBGDRH20, PZTC20, OGD23]. Varying [PLRR21]. Vendace [MVK23]. Video [HRSD21]. View [BHJ^{+}22]. Visible [HMJW20]. Vital [DYG^{+}23].

Walleye [CEL^{+}20, JZH^{+}23, RFR^{+}22a, RFR^{+}22b, VCV23]. Walleyes [JHM^{+}21]. Want [Nie21]. Warm [TDB20]. warmer [DGG^{+}23]. Water [AKB^{+}20, DKS^{+}22, EKO^{+}21, BHC^{+}23, EME23]. Waters [OGT22]. Watershed [JSI^{+}21, PO20, HCS23, MWS^{+}23, PM23]. watershed-level [HCS23]. Watershed-Scale [JSI^{+}21]. Weakfish [MHO^{+}22]. Welfare [RAZ^{+}21]. western [DYG^{+}23, PARM23, AKB^{+}20, CWLJ21]. Westslope [DEBW23]. Whelk [AFB^{+}22]. which [MVK23]. White [BRJ^{+}22, BSC^{+}21, OGT22]. Whitefish [DB23, Hax23, IBT^{+}20]. Whose [ALH21]. Wild [BKVH20, LCM^{+}21, MLZ^{+}20, TR20, LBC^{+}23]. Will [KGP22]. wind [BE23]. wind-driven [BE23]. Winter [DCB^{+}21, FJ20]. Wisconsin [IRE^{+}22, THH^{+}22]. Within [EECA20, DB23, NKL^{+}20]. Within-Season [EECA20]. Writings [Ken22].

Yakima [KSG^{+}22]. Year [MHO^{+}22]. Yearly [GYL^{+}21]. Years [PLRR21, TDB20, WCSJ20]. Yellow [AMZC^{+}20, BRJ^{+}22]. Yellowstone [HGW^{+}21, KTG^{+}20, MH20, VACS^{+}22]. York [BRJ^{+}22, GBR^{+}21, HWK20]. Young [MHO^{+}22, MVK23]. Young-of-the-Year [MHO^{+}22]. Yukon [EME23].

References

Anderson:2021:PSC

[ACAL21] Austin J. Anderson, Andrew M. Claiborne, Mickey Agha, and Marisa N. C. Litz. Puget Sound chum salmon growth linked to competitor abundance, climate indices, and copepod species richness. Transactions of the American Fisheries Society, 150 (6):707–729, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Archdeacon:2023:HSI

[ADR^{+}23] Thomas P. Archdeacon, Robert K. Dudley, W. Jason Remshardt, William Knight, Manuel Ulibarri, and Eric J. Gonzales. Hatchery supplementation increases potential spawning stock of Rio Grande Silvery Minnow after population bottlenecks. Transactions of the American Fisheries Society, 152 (2):187–200, March 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Askin:2022:PGS

[AFB^{+}22] Samantha E. Askin, Robert A. Fisher, Ellen E. Biesack, Rick Robins, and Jan R. McDowell. Population genetic structure in channeled whelk Busycotypus canaliculatus along the U.S. Atlantic Coast. Transactions of the American Fisheries Society, 151(5):543–558, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Aragon-Flores:2022:TER

[AFRHSB^{+}22] Edgar Abraham Aragón-Flores, Rocío Rodiles-Hernández, Miriam Soria-Barreto, Carmen G. Montaña, María Mercedes Castillo, and Matteo Cazzanelli. Trophic ecology and resource partitioning of common snook and Mexican snook in the Usumacinta River Delta. Transactions of the American Fisheries Society, 151(2):185–201, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Atkore:2020:MIW

[AKB^{+}20] Vidyadhar Atkore, Nachiket Kelkar, Shrinivas Badiger, Kartik Shanker, and Jagdish Krishnaswamy. Multiscale investigation of water chemistry effects on fish guild species richness in regulated and nonregulated rivers of India’s Western Ghats: Implications for restoration. Transactions of the American Fisheries Society, 149(3):298–319, May 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Arakawa:2021:WKD

[ALH21] Hiroaki Arakawa, Ralph T. Lampman, and Jon E. Hess. Whose kids did you eat? Genetic identification of species and parents of larval lampreys in fish predator guts. Transactions of the American Fisheries Society, 150(5):551–559, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:2020:EIM

[AMZC^{+}20] Brandon M. Armstrong, Francisco X. Mora-Zamorano, Michael J. Carvan III, Scott McNaught, Niladri Basu, Jessica Head, Rebekah H. Klingler, Lori Ivan, and Cheryl A. Murphy. Exploring the impacts of methylmercury-induced behavioral alterations in larval yellow perch in Lake Michigan using an individual-based model. Transactions of the American Fisheries Society, 149(6): 664–680, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2020:IIa

[Ano20a] Anonymous. Issue information. Transactions of the American Fisheries Society, 149(1):1–2, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2020:IIb

[Ano20b] Anonymous. Issue information. Transactions of the American Fisheries Society, 149(2):133–134, March 2020. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2020:IIc

[Ano20c] Anonymous. Issue information. Transactions of the American Fisheries Society, 149(3):243–244, May 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2020:IId

[Ano20d] Anonymous. Issue information. Transactions of the American Fisheries Society, 149(4):367–368, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2020:IIe

[Ano20e] Anonymous. Issue information. Transactions of the American Fisheries Society, 149(5):516, September 2020. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2020:IIf

[Ano20f] Anonymous. Issue information. Transactions of the American Fisheries Society, 149(6):649–650, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2020:RA

[Ano20g] Anonymous. Reviewer acknowledgments. Transactions of the American Fisheries Society, 149(6):770–771, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2021:IIa

[Ano21a] Anonymous. Issue information. Transactions of the American Fisheries Society, 150(1):1–2, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2021:IIb

[Ano21b] Anonymous. Issue information. Transactions of the American Fisheries Society, 150(2):145–146, March 2021. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2021:IIc

[Ano21c] Anonymous. Issue information. Transactions of the American Fisheries Society, 150(3):275–276, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2021:IId

[Ano21d] Anonymous. Issue information. Transactions of the American Fisheries Society, 150(4):423–424, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2021:IIe

[Ano21e] Anonymous. Issue information. Transactions of the American Fisheries Society, 150(5):549–550, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2021:IIf

[Ano21f] Anonymous. Issue information. Transactions of the American Fisheries Society, 150(6):665–666, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2021:RA

[Ano21g] Anonymous. Reviewer acknowledgments. Transactions of the American Fisheries Society, 150(6):807–808, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2022:E

[Ano22a] Anonymous. Erratum. Transactions of the American Fisheries Society, 151(2):260, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIa

[Ano22b] Anonymous. Issue information. Transactions of the American Fisheries Society, 151(1):1–2, January 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2022:IIb

[Ano22c] Anonymous. Issue information. Transactions of the American Fisheries Society, 151(2):125–126, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IIc

[Ano22d] Anonymous. Issue information. Transactions of the American Fisheries Society, 151(3):261–262, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Anonymous:2022:IId

[Ano22e] Anonymous. Issue information. Transactions of the American Fisheries Society, 151(4):397–398, July 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2022:IIe

[Ano22f] Anonymous. Issue information. Transactions of the American Fisheries Society, 151(5):525–526, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2022:IIf

[Ano22g] Anonymous. Issue information. Transactions of the American Fisheries Society, 151(6):655–656, November 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2023:IIa

[Ano23a] Anonymous. Issue information. Transactions of the American Fisheries Society, 152(1):1–2, January 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2023:IIb

[Ano23b] Anonymous. Issue information. Transactions of the American Fisheries Society, 152(2):127–128, March 2023. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2023:IIc

[Ano23c] Anonymous. Issue information. Transactions of the American Fisheries Society, 152(3):271–272, May 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2023:IId

[Ano23d] Anonymous. Issue information. Transactions of the American Fisheries Society, 152(4):381–382, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2023:IIe

[Ano23e] Anonymous. Issue information. Transactions of the American Fisheries Society, 152(5):513–514, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2023:IIf

[Ano23f] Anonymous. Issue information. Transactions of the American Fisheries Society, 152(6):709–710, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:2023:RA

[Ano23g] Anonymous. Reviewer acknowledgments. Transactions of the American Fisheries Society, 152(6):851–852, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashton:2021:ETF

[ARH^{+}21] Neil K. Ashton, Tyler J. Ross, Ryan S. Hardy, Sarah M. Stephenson, Valerie Evans, Nathan R. Jensen, Shawn P. Young, and Kenneth D. Cain. Effects of temperature fluctuation on burbot embryos: Implications of hydropower and climate change. Transactions of the American Fisheries Society, 150 (5):605–617, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brick:2022:RNS

[BAF22] J. Michael Brick, William R. Andrews, and John Foster. A review of nonprobability sampling using mobile apps for fishing effort and catch surveys. Transactions of the American Fisheries Society, 151(1):42–49, January 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bart:2021:CPG

[BDW21] Ryan J. Bart, Dennis R. DeVries, and Russell A. Wright. Change in piscivore growth potential after the introduction of a nonnative prey fish: a bioenergetics analysis. Transactions of the American Fisheries Society, 150(2):175–188, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bell:2023:TRM

[BE23] Geoffrey W. Bell and David B. Eggleston. Timing and route of migration of mature female blue crabs in a large, wind-driven estuary. Transactions of the American Fisheries Society, 152 (2):253–269, March 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bzonek:2022:DMI

[BEH^{+}22] P. A. Bzonek, P. D. Edwards, C. T. Hasler, C. D. Suski, R. Boonstra, and N. E. Mandrak. Deterring the movement of an invasive fish: Individual variation in common carp responses to acoustic and stroboscopic stimuli. Transactions of the American Fisheries Society, 151(1):112–123, January 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burns:2023:EMP

[BGR^{+}23] Christopher W. Burns, Pascale Gibeau, David Robichaud, Cameron McCulloch, Julio Novoa, and Krystal Lockert. Eulachon migration patterns in the lower Fraser River revealed through acoustic telemetry. Transactions of the American Fisheries Society, 152(5):632–649, September 2023. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burdick:2020:MEJ

[BHC^{+}20] Summer M. Burdick, Danielle M. Hereford, Carla M. Conway, Nathan V. Banet, Rachel Powers, Barbara A. Martin, and Diane G. Elliott. Mortality of endangered juvenile Lost River suckers associated with cyanobacteria blooms in mesocosms in Upper Klamath Lake, Oregon. Transactions of the American Fisheries Society, 149(3):245–265, May 2020. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Booth:2023:EMC

[BHC^{+}23] Ian Taylor Booth, Kyle J. Hartman, Derek Crane, Jeff Hansbarger, Jordan Weeks, Josh Henesy, Heather Walsh, and Jeff Williams. Evaluating muskellunge catch-and-release mortality at elevated summer water temperature. Transactions of the American Fisheries Society, 152(5):577–593, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Breece:2021:SIT

[BHF21] M. W. Breece, A. L. Higgs, and D. A. Fox. Spawning intervals, timing, and riverine habitat use of adult Atlantic sturgeon in the Hudson River. Transactions of the American Fisheries Society, 150(4):528–537, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boughton:2022:CTS

[BHJ^{+}22] David A. Boughton, Lee R. Harrison, Sara N. John, Rosealea M. Bond, Colin L. Nicol, Carl J. Legleiter, and Ryan T. Richardson. Capacity of two Sierra Nevada rivers for reintroduction of anadromous salmonids: Insights from a high-resolution view. Transactions of the American Fisheries Society, 151(1):13–41, January 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budy:2021:ECR

[BHTH21] Phaedra Budy, Nicholas A. Heredia, Gary P. Thiede, and Erik Horgen. Exploring the contemporary relationship between predator and prey in a significant, reintroduced Lahontan cutthroat trout population. Transactions of the American Fisheries Society, 150(3):291–306, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burgad:2023:LTP

[BKM23] Aaron A. Burgad, Brian R. Kesner, and Paul C. Marsh. Longterm patterns of fish community structure and decline in native species in a Colorado River reservoir, Arizona and Nevada.

Transactions of the American Fisheries Society, 152(6):760– 771, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bruce:2020:SFI

[BKVH20] Spencer A. Bruce, Yuka Kutsumi, Chris Van Maaren, and Matthew P. Hare. Stocked-fish introgression into wild brook trout populations depends on habitat. Transactions of the American Fisheries Society, 149(4):427–442, July 2020. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown-Peterson:2021:IDA

[BPLL21] Nancy J. Brown-Peterson, Robert T. Leaf, and Andrea J. Leontiou. Importance of depth and artificial structure as predictors of female red snapper reproductive parameters. Transactions of the American Fisheries Society, 150(1):115–129, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bradford:2020:EFA

[Bra20] Mike Bradford. Environmental flow assessment: Methods and applications. Transactions of the American Fisheries Society, 149(3):366, May 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brooking:2022:ERG

[BRJ^{+}22] Thomas E. Brooking, Lars G. Rudstam, James R. Jackson, Anthony J. VanDeValk, Kristen T. Holeck, Christopher W. Hotaling, and John E. Cooper. Effects of round goby on the benthic invertebrate community and on diets and growth of yellow perch and white perch in Oneida Lake, New York. Transactions of the American Fisheries Society, 151(5):641–654, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budnik:2021:EHH

[BSC^{+}21] Richard R. Budnik, Geoffrey B. Steinhart, Joseph D. Conroy, Rebecca A. Dillon, Richard D. Zweifel, and Stuart A. Ludsin. Effects of hypoxia on habitat quality of reservoir largemouth bass, saugeye, and white crappie. Transactions of the American Fisheries Society, 150(1):75–88, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brodie:2022:DSM

[BSP^{+}22] Russell B. Brodie, Justin J. Solomon, Richard Paperno, Theodore S. Switzer, and Christopher E. Swanson. Development of survey methods to effectively sample juvenile red snapper along the Atlantic Coast of Florida. Transactions of the American Fisheries Society, 151(5):571–586, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bradford:2023:DFL

[BT23] Michael J. Bradford and Garth C. Taylor. Diversity in freshwater life history in spring and summer Chinook salmon from the Fraser River, Canada. Transactions of the American Fisheries Society, 152(2):129–144, March 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budy:2020:MRP

[BTM^{+}20] Phaedra Budy, Paul D. Thompson, Matt D. McKell, Gary P. Thiede, Tim E. Walsworth, and Mary M. Conner. A multifaceted reconstruction of the population structure and life history expressions of a remnant metapopulation of Bonneville cutthroat trout: Implications for maintaining intermittent connectivity. Transactions of the American Fisheries Society, 149(4): 443–461, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bohling:2020:GCC

[BVB20] Justin Bohling, Jennifer Von Bargen, and Peter Bahls. Genomic characterization of Coho salmon spawning populations from the Hood Canal. Transactions of the American Fisheries Society, 149(1):3–13, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Booher:2022:ITS

[BW22] Evan C. J. Booher and Annika W. Walters. Identifying translocation sites for a climate relict population of finescale dace. Transactions of the American Fisheries Society, 151(2):245– 259, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Campbell:2023:MPP

[CBB^{+}23] Matthew A. Campbell, Mary E. Badger, Nick Buckmaster, Andrew B. Starostka, Travis Hawks, and Amanda J. Finger. Molecular phylogenetic and population genetic relationships of a putative species of sucker (Catostomus sp.) from Surprise Valley in the Great Basin, USA. Transactions of the American Fisheries Society, 152(3):273–286, May 2023. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ciancio:2020:EDP

[CBF^{+}20] Javier Ernesto Ciancio, Saia Bartes, Santiago Fernández, Camila Harillo, and Julio Lancelotti. Energy density predictors for Argentine anchovy Engraulis anchoita, a key species of the Southwestern Atlantic Ocean. Transactions of the American Fisheries Society, 149(2):204–212, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chen:2020:RSR

[CEL^{+}20] Kuan-Yu Chen, Peter T. Euclide, Stuart A. Ludsin, Wesley A. Larson, Michael G. Sovic, H. Lisle Gibbs, and Elizabeth A. Marschall. RAD-Seq refines previous estimates of genetic structure in Lake Erie walleye. Transactions of the American Fisheries Society, 149(2):159–173, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coen:2021:HLG

[CFL^{+}21] Amanda E. Coen, Max Fish, Randy Lovell, Jeff Rodzen, Rachel Schwartz, and Andrea Schreier. High levels of genetic divergence detected in Sacramento perch suggests two divergent translocation sources. Transactions of the American Fisheries Society, 150(3):375–387, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colvin:2020:EFA

[CHDF20] S. A. R. Colvin, B. S. Helms, D. R. DeVries, and J. W. Feminella. Environmental and fish assemblage contrasts in blackwater and clearwater streams. Transactions of the American Fisheries Society, 149(3):335–349, May 2020. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crossman:2022:PRS

[CWKY22] James A. Crossman, Molly A. H. Webb, Josh Korman, and Michael D. Yard. Population reproductive structure of rainbow trout determined by histology and advancing methods to assign sex and assess spawning capability. Transactions of the American Fisheries Society, 151(4):422–440, July 2022. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:2021:ATA

[CWLJ21] Benjamin J. Clemens, Matthew A. Weeber, Mark Lewis, and Michelle Jones. Abundance trends for adult Pacific lamprey in Western Oregon (USA): Historic declines, recent increases, and relative contributions from coastal rivers. Transactions of the American Fisheries Society, 150(6):761–776, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drew:2023:OGT

[DB23] Katie A. Drew and Randy J. Brown. Opposing growth trends in juvenile and mature humpback whitefish within the upper Tanana River drainage, Alaska, revealed using otolith biochronology. Transactions of the American Fisheries Society, 152(4):415–428, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dale:2021:IIG

[DCB^{+}21] Leah L. Dale, James Patrick Cronin, Virginia L. Brink, Blair E. Tirpak, John M. Tirpak, and William E. Pine III. Identifying information gaps in predicting winter foraging habitat for juvenile Gulf sturgeon. Transactions of the American Fisheries Society, 150(2):222–241, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Denison:2023:SMC

[DCF^{+}23] Colby D. Denison, Amy Cottrell, Troy M. Farmer, Dewayne A. Fox, David M. Hood, William C. Post, Gregory Sorg, Ellen Waldrop, and Brandon K. Peoples. Seasonal migration cues differ for dual-spawning Atlantic sturgeon in the Great Pee Dee River. Transactions of the American Fisheries Society, 152(5): 694–708, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dangora:2023:EES

[DEBW23] Anthony J. Dangora, Lisa A. Eby, Craig A. Barfoot, and Andrew R. Whiteley. Evaluating the effects of selective passage of migratory westslope cutthroat trout on nonnative admixture. Transactions of the American Fisheries Society, 152(6): 772–787, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doll:2020:LTP

[DERF20] Jason C. Doll, Luke Etchison, William T. Russ, and Stephen J. Fraley. Long-term population dynamics and habitat association of the federally threatened spotfin chub in the Little Tennessee River. Transactions of the American Fisheries Society, 149(5): 587–599, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dauwalter:2023:RAT

[DGG^{+}23] Daniel C. Dauwalter, Rosalinda Gonzalez, Tim Gatewood, Zachary S. Beard, and Zachary Jackson. Resiliency of Apache Trout habitats in a warmer and drier climate future. Transactions of the American Fisheries Society, 152(3):361–380, May 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dula:2022:EHM

[DKD^{+}22] Brendan T. Dula, Adam J. Kaeser, Mark J. D’Ercole, Cecil A. Jennings, and Adam G. Fox. Effects of Hurricane Michael on Gulf sturgeon in the Apalachicola River system, Florida. Transactions of the American Fisheries Society, 151(6):725– 742, November 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dinken:2022:EWT

[DKS^{+}22] Colin P. Dinken, Kevin R. Keretz, Harold L. Schramm, Jr., Lora Petrie-Hanson, M. Wes Schilling, and Peter J. Allen. The effects of water temperature and simulated angling on the physiological stress response of largemouth bass. Transactions of the American Fisheries Society, 151(4):487–506, July 2022. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Devine:2021:FSA

[DRR^{+}21] Matthew T. Devine, Julianne Rosset, Allison H. Roy, Benjamin I. Gahagan, Michael P. Armstrong, Andrew R. Whiteley, and Adrian Jordaan. Feeling the squeeze: Adult run size and habitat availability limit juvenile river herring densities in lakes. Transactions of the American Fisheries Society, 150 (2):207–221, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dunnigan:2021:FIR

[DT21] James L. Dunnigan and Marc M. Terrazas. Factors influencing rainbow trout annual growth in a large regulated river. Transactions of the American Fisheries Society, 150(1):89–114, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dzul:2023:VRB

[DYG^{+}23] Maria Dzul, Charles B. Yackulic, Mariah Giardina, David R. Van Haverbeke, and Michael Yard. Vital rates of a burgeoning population of humpback chub in western Grand Canyon. Transactions of the American Fisheries Society, 152(4):443–459, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Enterline:2020:QWS

[EECA20] Claire L. Enterline, Scott P. Elzey, Bradford C. Chase, and Matthew H. Ayer. Quantifying within-season repeat spawning by rainbow smelt with implications for comparing sex ratios and survey catches among runs. Transactions of the American Fisheries Society, 149(1):27–42, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eshenroder:2020:SCE

[EJ20] Randy L. Eshenroder and Peter C. Jacobson. Speciation in cisco with emphasis on secondary contacts, plasticity, and hybridization. Transactions of the American Fisheries Society, 149(6): 721–740, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eshenroder:2021:RTA

[EKO^{+}21] Randy L. Eshenroder, Yu-Chun Kao, Timothy P. O’Brien, Chris M. Olds, Chris L. Davis, and Alexander T. Duncan. Replacement of the typical artedi form of Cisco Coregonus artedi in Lake Huron by endemic shallow-water ciscoes, including putative hybrids. Transactions of the American Fisheries Society, 150(6):792–806, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eiler:2023:SDW

[EME23] John H. Eiler, Michele M. Masuda, and Allison N. Evans. Swimming depths and water temperatures encountered by radioarchival-tagged Chinook salmon during their spawning migration in the Yukon River basin. Transactions of the American Fisheries Society, 152(1):51–74, January 2023. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erdman:2022:BPS

[EMG^{+}22] Bradley Erdman, Matthew G. Mitro, Joanna D. T. Griffin, David Rowe, David C. Kazyak, Keith Turnquist, Michael Siepker, Loren Miller, Wendylee Stott, Michael Hughes, Brian Sloss, Michael T. Kinnison, Chris C. Wilson, and Wes Larson. Broadscale population structure and hatchery introgression of Midwestern brook trout. Transactions of the American Fisheries Society, 151(1):81–99, January 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Euclide:2020:GSS

[ERH^{+}20] Peter T. Euclide, Jenna Ruzich, Scott P. Hansen, David Rowe, Troy G. Zorn, and Wesley A. Larson. Genetic structure of smallmouth bass in the Lake Michigan and Upper Mississippi River drainages relates to habitat, distance, and drainage boundaries. Transactions of the American Fisheries Society, 149(4):383– 397, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Funnell:2023:SSE

[FBK^{+}23] Tyler R. Funnell, Travis O. Brenden, Richard Kraus, Tom Mac-Dougall, James Markham, Charles Murray, Jason Robinson, and Christopher S. Vandergoot. Seasonal spatial ecology of lake trout in Lake Erie. Transactions of the American Fisheries Society, 152(5):672–693, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Faulkner:2021:AAF

[FBW^{+}21] James R. Faulkner, Blane L. Bellerud, Daniel L. Widener, Steven G. Smith, and Richard W. Zabel. Associations among fish length, dam passage history, and survival to adulthood in two at-risk species of Pacific salmon: Response to comment. Transactions of the American Fisheries Society, 150(2):196– 206, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See [SHS^{+}21].

Frechette:2021:ASM

[FDDB21] Danielle M. Frechette, Mélanie Dionne, Julian J. Dodson, and Normand E. Bergeron. Atlantic salmon movement patterns and habitat use during colonization of novel habitat. Transactions of the American Fisheries Society, 150(3):327–344, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:2020:OST

[FDS^{+}20] Gregory S. Fraser, Patrick W. DeHaan, Christian T. Smith, Jennifer F. Von Bargen, Mathew R. Cooper, and Thomas J. Desgrosseillier. Overlap of spatial and temporal spawning distributions of spring and summer Chinook salmon results in hybridization in the Upper Columbia River. Transactions of the American Fisheries Society, 149(5):517–531, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Favrot:2020:FWM

[FJ20] Scott D. Favrot and Brian C. Jonasson. Fall and winter movement dynamics of naturally produced spring Chinook salmon parr in two neighboring interior Pacific Northwest natal rivers. Transactions of the American Fisheries Society, 149(5):532– 551, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fredrickson:2022:NFC

[FRF^{+}22] Anthea Fredrickson, Guillaume Rieucau, Quenton Fontenot, Alec Lackmann, and Solomon R. David. Nonlethal fin clip model validation for stable isotope analysis of spotted gar and alligator gar. Transactions of the American Fisheries Society, 151(1):72–80, January 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

George:2021:EER

[GBR^{+}21] Scott D. George, Barry P. Baldigo, Christopher B. Rees, Meredith L. Bartron, and Dylan Winterhalter. Eastward expansion of round goby in New York: Assessment of detection methods and current range. Transactions of the American Fisheries Society, 150(2):258–273, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

George:2023:UED

[GBR^{+}23] Scott D. George, Barry P. Baldigo, Christopher B. Rees, Meredith L. Bartron, John J. Wiley Jr, Daniel S. Stich, Scott M. Wells, and Dylan R. Winterhalter. Use of environmental DNA to assess American eel distribution, abundance, and barriers in a river–canal system. Transactions of the American Fisheries Society, 152(3):310–326, May 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gilligan-Lunda:2021:CCM

[GLSM^{+}21] Erin K. Gilligan-Lunda, Daniel S. Stich, Katherine E. Mills, Michael M. Bailey, and Joseph D. Zydlewski. Climate change may cause shifts in growth and instantaneous natural mortality of American shad throughout their native range. Transactions of the American Fisheries Society, 150(3):407–421, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galbreath:2022:PMH

[GSN^{+}22] Peter F. Galbreath, Benjamin A. Staton, Hayley M. Nuetzel, Chad A. Stockton, Curtis M. Knudsen, Lea R. Medeiros, Ilana J. Koch, William J. Bosch, and Andrew L. Pierce. Precocious maturation of hatchery-raised spring Chinook salmon as age-2 minijacks is not detectably affected by sire age. Transactions of the American Fisheries Society, 151(3):333–346, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Ghinter:2021:LSY

[GYL^{+}21] Léopold Ghinter, Wahiba Ait Youcef, Yvan Lambert, M. Joanne Morgan, and Céline Audet. Local, seasonal, and yearly condition of juvenile Greenland halibut revealed by the Le Cren Condition Index. Transactions of the American Fisheries Society, 150(6):730–747, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haxton:2023:BCI

[Hax23] Tim Haxton. Biological characteristics of inland lake whitefish populations in Ontario. Transactions of the American Fisheries Society, 152(6):819–834, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hughes:2020:CFI

[HBH^{+}20] Robert M. Hughes, George Boxall, Alan T. Herlihy, Jason Adams, and Daniel B. Young. A complete fisheries inventory of the Chulitna River Basin, Lake Clark National Park and Preserve, Alaska: Example of a minimally disturbed basin. Transactions of the American Fisheries Society, 149(1):14–26, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herigan:2023:AEL

[HCS23] Garrett M. Herigan, Derek P. Crane, and Mark C. Scott. Association of an endemic leuciscid, the Sandhills Chub, with microhabitat features and watershed-level habitat characteristics. Transactions of the American Fisheries Society, 152(2):217– 231, March 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hess:2022:PLT

[HDJ^{+}22] Jon E. Hess, Thomas A. Delomas, Aaron D. Jackson, Michael J. Kosinski, Mary L. Moser, Laurie L. Porter, Greg Silver, Tod Sween, Laurie A. Weitkamp, and Shawn R. Narum. Pacific lamprey translocations to the Snake River boost abundance of all life stages. Transactions of the American Fisheries Society, 151(3):263–296, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Hansell:2020:IHK

[HDKC20] Alexander C. Hansell, Gregory R. DeCelles, Michael E. Kersula, and Steven X. Cadrin. Incorporating harvesters’ knowledge into an index of abundance for Atlantic halibut in the Northwest Atlantic. Transactions of the American Fisheries Society, 149 (6):741–752, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hershey:2022:EFP

[HDW^{+}22] Henry Hershey, Dennis R. DeVries, Russell A. Wright, Dustin McKee, and David L. Smith. Evaluating fish passage and tailrace space use at a low-use low-head lock and dam. Transactions of the American Fisheries Society, 151(1):50–71, January 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heredia:2021:ILT

[HGW^{+}21] Nicholas A. Heredia, Robert E. Gresswell, Molly A. H. Webb, Travis O. Brenden, and Philip T. Sandstrom. Invasive lake trout reproduction in Yellowstone Lake under an active suppression program. Transactions of the American Fisheries Society, 150 (5):637–650, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hanlon:2023:IHO

[HJ23] Charles G. Hanlon and Alyssa Jordan. Influence of habitat and other factors on largemouth bass abundance in Lake Okeechobee, Florida. Transactions of the American Fisheries Society, 152(4):429–442, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Huntsman:2021:IED

[HLC21] Brock M. Huntsman, Abigail J. Lynch, and Colleen A. Caldwell. Interacting effects of density-dependent and densityindependent factors on growth rates in southwestern cutthroat trout populations. Transactions of the American Fisheries Society, 150(5):651–664, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harstad:2023:ERD

[HLC^{+}23] Deborah L. Harstad, Donald A. Larsen, Lance Clarke, Dina K. Spangenberg, Robert Hogg, Brett Requa, and Brian R. Beckman. The effect of reducing dietary lipid and food availability on precocious male maturation in Chinook salmon: a productionscale hatchery experiment. Transactions of the American Fisheries Society, 152(2):232–252, March 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haro:2020:EVL

[HMJW20] Alex Haro, Scott Miehls, Nicholas S. Johnson, and C. Michael Wagner. Evaluation of visible light as a cue for guiding downstream migrant juvenile sea lamprey. Transactions of the American Fisheries Society, 149(5):635–647, September 2020. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Helminen:2021:MTF

[HOL21] Jani Helminen, Antóin M. O’Sullivan, and Tommi Linnansaari. Measuring tailbeat frequencies of three fish species from adaptive resolution imaging sonar data. Transactions of the American Fisheries Society, 150(5):627–636, September 2021. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hess:2021:RRP

[HPM^{+}21] Jon E. Hess, Rebecca L. Paradis, Mary L. Moser, Laurie A. Weitkamp, Thomas A. Delomas, and Shawn R. Narum. Robust recolonization of Pacific lamprey following dam removals. Transactions of the American Fisheries Society, 150(1):56–74, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunter:2020:GFR

[HRS^{+}20] Robert D. Hunter, Edward F. Roseman, Nicholas M. Sard, Robin L. DeBruyne, Jinliang Wang, and Kim T. Scribner. Genetic family reconstruction characterizes lake sturgeon use of newly constructed spawning habitat and larval dispersal. Transactions of the American Fisheries Society, 149(3):266–283, May 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hitt:2021:CUV

[HRSD21] Nathaniel P. Hitt, Karli M. Rogers, Craig D. Snyder, and C. Andrew Dolloff. Comparison of underwater video with electrofishing and dive counts for stream fish abundance estimation. Transactions of the American Fisheries Society, 150(1):24–37, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holcomb:2020:USH

[HSJ^{+}20] Kathryn M. Holcomb, Paul Schueller, Howard L. Jelks, John R. Knight, and Micheal S. Allen. Use of strong habitat–abundance relationships in assessing population status of cryptic fishes: an example using the harlequin darter. Transactions of the American Fisheries Society, 149(3):320–334, May 2020. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hager:2020:SDF

[HWK20] Christian H. Hager, J. Carter Watterson, and Jason E. Kahn. Spawning drivers and frequency of endangered Atlantic sturgeon in the York River system. Transactions of the American Fisheries Society, 149(4):474–485, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isermann:2020:DNG

[IBT^{+}20] Daniel A. Isermann, Matthew J. Belnap, Keith N. Turnquist, Brian L. Sloss, Justin A. VanDeHey, Scott P. Hansen, and David C. Caroffino. Defining the need for genetic stock assignment when describing stock demographics and dynamics: an example using lake whitefish in Lake Michigan. Transactions of the American Fisheries Society, 149(4):398–413, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ivanova:2021:MEP

[IJF21] Silviya V. Ivanova, Timothy B. Johnson, and Aaron T. Fisk. Movement ecology of a potamodromous top predator in a large lake: Synchrony and coexistence of distinct migratory patterns. Transactions of the American Fisheries Society, 150(6): 748–760, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isermann:2022:LSM

[IRE^{+}22] Daniel A. Isermann, Joshua K. Raabe, Emma G. Easterly, Joshua C. Schulze, Nicholas J. Porter, Daniel J. Dembkowski, Michael C. Donofrio, Darren R. Kramer, and Robert F. Elliott. Lake sturgeon movement after trap and transfer around two dams on the Menominee River, Wisconsin–Michigan. Transactions of the American Fisheries Society, 151(5):611–629, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Izzo:2022:SMS

[IZMP22] Lisa K. Izzo, Gayle Barbin Zydlewski, J. Ellen Marsden, and Donna L. Parrish. Seasonal movements and spatial overlap of juvenile and adult lake sturgeon in Lake Champlain. Transactions of the American Fisheries Society, 151(6):666–681, November 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Janzen:2023:EGD

[JBD23] Francesco H. Janzen and Gabriel Blouin-Demers. Estimating the genetic diversity and potential influence of habitat segregation in channel catfish. Transactions of the American Fisheries Society, 152(6):726–737, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2023:BMN

[JBO23] Rachelle C. Johnson, David A. Beauchamp, and Julian D. Olden. Bioenergetics model for the nonnative Redside Shiner. Transactions of the American Fisheries Society, 152(1):94–113, January 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2021:GST

[JDZ^{+}21] Garret R. Johnson, Rebecca A. Dillon, Richard D. Zweifel, Stuart A. Ludsin, and Joseph D. Conroy. Gizzard shad target strength-to-body size equations at multiple hydroacoustic frequencies. Transactions of the American Fisheries Society, 150 (2):242–257, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnston:2021:MEE

[JHM^{+}21] T. A. Johnston, A. J. Harry, R. D. Montgomerie, M. D. Wiegand, G. A. Spiers, J. M. Casselman, and W. C. Leggett. Maternal effects on embryonic development and survival in walleyes of Lake Nipissing, Ontario. Transactions of the American Fisheries Society, 150(6):777–791, November 2021. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jenkins:2020:EPC

[JPC^{+}20] Laura E. Jenkins, Andrew L. Pierce, Christopher C. Caudill, Neil D. Graham, Lea R. Medeiros, Douglas R. Hatch, and James J. Nagler. Effects of physiological condition on aspects of repeat spawning in female steelhead reconditioned in captivity. Transactions of the American Fisheries Society, 149(2):213– 224, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Joy:2021:JCC

[JSI^{+}21] Philip J. Joy, Craig A. Stricker, Renae Ivanoff, Shiao Y. Wang, Mark S. Wipfli, Andrew C. Seitz, Jiaqi Huang, and Mathew B. Tyers. Juvenile Coho and Chinook salmon growth, size, and condition linked to watershed-scale salmon spawner abundance. Transactions of the American Fisheries Society, 150(3):307– 326, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jacobs:2021:CFR

[JTB^{+}21] Gregory R. Jacobs, Russell F. Thurow, John M. Buffington, Daniel J. Isaak, and Seth J. Wenger. Climate, fire regime, geomorphology, and conspecifics influence the spatial distribution of Chinook salmon redds. Transactions of the American Fisheries Society, 150(1):8–23, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See erratum [JTB^{+}22].

Jacobs:2022:ECF

[JTB^{+}22] Gregory R. Jacobs, Russell F. Thurow, John M. Buffington, Daniel J. Isaak, and Seth J. Wenger. Erratum: Climate, fire regime, geomorphology, and conspecifics influence the spatial distribution of Chinook salmon redds. Transactions of the American Fisheries Society, 151(3):390–395, May 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic). See [JTB^{+}21].

Jacobson:2020:MVC

[JZGE20] Peter C. Jacobson, Kyle D. Zimmer, Ryan Grow, and Randy L. Eshenroder. Morphological variation of cisco across gradients of lake productivity. Transactions of the American Fisheries Society, 149(4):462–473, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:2023:GEN

[JZH^{+}23] Andrew Johnson, Katherine Zipfel, Eric Hallerman, Wade Massure, Peter Euclide, and Amy Welsh. Genomic evaluation of native walleye in the Appalachian region and the effects of stocking. Transactions of the American Fisheries Society, 152(3): 346–360, May 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kratina:2023:UFH

[KDW^{+}23] Garret J. Kratina, Dennis R. DeVries, Russell A. Wright, Eric Peatman, Steven J. Rider, and Honggang Zhao. Using fish hard-part microchemistry and genetics to quantify population impacts of low-use lock-and-dam structures on the Alabama River. Transactions of the American Fisheries Society, 152(4): 490–512, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kenchington:2022:SWS

[Ken22] Trevor Kenchington. Selected writings of Sidney J. Holt: Documenting a lifelong pursuit of sustainable and profitable fishing. Transactions of the American Fisheries Society, 151(3): 389, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Krolow:2022:IYB

[KGP22] Aaron D. Krolow, Aaron D. Geheber, and Kyle R. Piller. If you build it, will they come? An environmental DNA assessment of fish assemblages on artificial reefs in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 151 (3):297–321, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Kemp:2023:FFR

[KK23] Stanley J. Kemp and Mary J. Kemp. A flooded future for river chub? Future impacts of climate change and urbanization on reproduction of a keystone native fish species. Transactions of the American Fisheries Society, 152(5):594–609, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kazyak:2022:PGB

[KLK^{+}22] David C. Kazyak, Barbara A. Lubinski, Matt A. Kulp, Kasey C. Pregler, Andrew R. Whiteley, Eric Hallerman, Jason A. Coombs, Yoichiro Kanno, Jacob M. Rash, Raymond P. Morgan II, Jim Habera, Jason Henegar, T. Casey Weathers, Matthew T. Sell, Anthony Rabern, Dan Rankin, and Tim L. King. Population genetics of brook trout in the Southern Appalachian Mountains. Transactions of the American Fisheries Society, 151(2): 127–149, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Koch:2022:ESU

[KSG^{+}22] Ilana J. Koch, Todd R. Seamons, Peter F. Galbreath, Hayley M. Nuetzel, Andrew P. Matala, Kenneth I. Warheit, David E. Fast, Mark V. Johnston, Charles R. Strom, Shawn R. Narum, and William J. Bosch. Effects of supplementation in Upper Yakima River Chinook salmon. Transactions of the American Fisheries Society, 151(3):373–388, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Kennedy:2022:SDP

[KSH^{+}22] Benjamen M. Kennedy, Matt J. Smith, John S. A. Holmes, Roger P. Root, Christian T. Smith, and Douglas P. Peterson. Size and distribution of parr produced from natural- and hatchery-origin steelhead spawning naturally in a small Pacific Northwest coastal stream. Transactions of the American Fisheries Society, 151(4):507–524, July 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelly:2021:FAD

[KSW21] Brett Kelly, Michael J. Siepker, and Michael J. Weber. Factors associated with detection and distribution of native brook trout and introduced brown trout in the Driftless Area of Iowa. Transactions of the American Fisheries Society, 150(3):388–406, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Koel:2020:OPD

[KTG^{+}20] Todd M. Koel, Nathan A. Thomas, Christopher S. Guy, Philip D. Doepke, Drew J. MacDonald, Alex S. Poole, Wendy M. Sealey, and Alexander V. Zale. Organic pellet decomposition induces mortality of lake trout embryos in Yellowstone Lake. Transactions of the American Fisheries Society, 149(1):57–70, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:2023:MIE

[KTJ^{+}23] Aron D. Katz, Sasha J. Tetzlaff, Mark D. Johnson, John D. Noble, Steven Rood, Derek Maki, and Jinelle H. Sperry. Molecular identification and environmental DNA detection of gill lice ectoparasites associated with brook trout declines. Transactions of the American Fisheries Society, 152(6):788–808, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LeCheminant:2021:MDS

[LBAW21] Alex G. LeCheminant, Gabriel M. Barrile, Shannon E. Albeke, and Annika W. Walters. Movement dynamics and survival of stocked Colorado River cutthroat trout. Transactions of the American Fisheries Society, 150(6):679–693, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lingard:2023:EIE

[LBC^{+}23] Stephanie A. Lingard, Arthur L. Bass, Katrina V. Cook, Michelle Fortier, Geoffrey G. Price, and Scott G. Hinch. Evaluating the influence of environmental and biological factors on migration behavior and residence duration of wild subyearling Chinook salmon in a fjord estuary using miniature acoustic transmitters. Transactions of the American Fisheries Society, 152(5):610–631, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lubich:2021:EFP

[LCFSS21] Chiara Lubich, Caroline Campos, Carlos Freitas, and Flávia Siqueira-Souza. Effects of fishing on the population of speckled pavon Cichla temensis in the Middle Negro River (Amazonas State, Brazil): A decrease in the size of the trophy fish? Transactions of the American Fisheries Society, 150(6): 667–678, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Losee:2021:MSP

[LCM^{+}21] James P. Losee, Andrew M. Claiborne, Gabe M. Madel, Matt Klungle, and Lance Campbell. Is marine survival for Puget Sound’s wild steelhead really that bad? A Nisqually River case study evaluating estimates of productivity and survival of Oncorhynchus mykiss. Transactions of the American Fisheries Society, 150(2):160–174, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lutz-Carrillo:2023:LHE

[LCSNB23] Dijar J. Lutz-Carrillo, Warren Schlechte, Jake Norman, and Daniel L. Bennett. Lineage and hybridization effects on size potential in the largemouth bass complex. Transactions of the American Fisheries Society, 152(2):145–168, March 2023. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laubach:2023:ETA

[LDWS23] Colin Laubach, Dennis R. DeVries, Russell A. Wright, and David L. Smith. The effect of temperature and acute hypoxia/ hyperoxia exposure on swimming performance and kinematics of freshwater drum. Transactions of the American Fisheries Society, 152(4):460–474, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

LovellFord:2020:PRD

[LFL^{+}20] Rachel M. LovellFord, Rebecca L. Flitcroft, Sarah L. Lewis, Mary V. Santelmann, and Gordon E. Grant. Patterns of river discharge and temperature differentially influence migration and spawn timing for Coho salmon in the Umpqua River Basin, Oregon. Transactions of the American Fisheries Society, 149 (6):695–708, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lin:2020:AGD

[LJ20] Yu-Jia Lin and Brian M. Jessop. Application of generalized depletion model to recruitment of American eel elvers and empirical support from survey data. Transactions of the American Fisheries Society, 149(5):576–586, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lawrence:2022:CRI

[LJC^{+}22] Michael J. Lawrence, Ken M. Jeffries, Steven J. Cooke, Eva C. Enders, Caleb T. Hasler, Christopher M. Somers, Cory D. Suski, and Michael J. Louison. Catch-and-release ice fishing: Status, issues, and research needs. Transactions of the American Fisheries Society, 151(3):322–332, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Legett:2021:DPR

[LJR^{+}21] Henry D. Legett, Adrian Jordaan, Allison H. Roy, John J. Sheppard, Marcelo Somos-Valenzuela, and Michelle D. Staudinger. Daily patterns of river herring (Alosa spp.) spawning migrations: Environmental drivers and variation among coastal streams in Massachusetts. Transactions of the American Fisheries Society, 150(4):501–513, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laiveling:2023:RCI

[LLB23] Audrey R. Laiveling, Christopher N. Lorentz, and Michael T. Booth. River connectivity increases the diversity of fish communities in gravel pit lakes. Transactions of the American Fisheries Society, 152(5):550–576, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Love:2021:UPI

[LN21] Joseph W. Love and Joshua J. Newhard. Using published information to predict consumption by northern snakehead in Maryland. Transactions of the American Fisheries Society, 150(4): 425–434, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lankford:2021:OCB

[LT21] Thomas E. Lankford, Jr. and Timothy E. Targett. Ontogenetic changes in behavioral and energetic responses of juvenile Atlantic croaker to an estuarine salinity gradient. Transactions of the American Fisheries Society, 150(5):618–626, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maynard:2020:MLC

[MB20] George A. Maynard and Zofia Baumann. Methylmercury levels in commercially harvested spiny dogfish captured off the coast of Massachusetts. Transactions of the American Fisheries Society, 149(4):486–497, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marcek:2020:IEH

[MBD^{+}20] Benjamin J. Marcek, Emily A. Burbacher, Konrad Dabrowski, Kim P. Winslow, and Stuart A. Ludsin. Interactive effects of hypoxia and temperature on consumption, growth, and condition of juvenile hybrid striped bass. Transactions of the American Fisheries Society, 149(1):71–83, January 2020. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCormick:2020:UIP

[MH20] Joshua L. McCormick and Brett High. Using an integrated population model to evaluate Yellowstone cutthroat trout responses to management actions. Transactions of the American Fisheries Society, 149(2):135–146, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCormick:2023:USS

[MH23] Joshua L. McCormick and John W. Heckel IV. Using statespace models to estimate recreational angling effort and infer processes that regulate angler dynamics. Transactions of the American Fisheries Society, 152(6):738–748, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McGarvey:2020:GAN

[MHI^{+}20] Lauren M. McGarvey, Leif J. Halvorson, Jason E. Ilgen, Christopher S. Guy, Jason G. McLellan, and Molly A. H. Webb. Gametogenesis and assessment of nonlethal tools to assign sex and reproductive condition in burbot. Transactions of the American Fisheries Society, 149(2):225–240, March 2020. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Michels:2021:CMF

[MHM21] N. O. Michels, T. R. Hrabik, and A. F. Mensinger. Comparison of multipiscivore foraging success on native and invasive prey fish under variable light intensities. Transactions of the American Fisheries Society, 150(3):361–374, May 2021. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miyamoto:2022:CDC

[MHN^{+}22] Kouta Miyamoto, Hiroki Hoshikawa, Jun Nagura, Yoshihiro Sawamoto, Miloš – Dor– dević, and Kengo Ohta. Cochineal dye concentration and treatment time for otolith marking of Japanese smelt Hypomesus nipponensis embryos. Transactions of the American Fisheries Society, 151(6):657–665, November 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mathews:2022:ETA

[MHO^{+}22] Joseph A. Mathews, Edward A. Hale, Haley A. Oleynik, John H. Clark, Ian A. Park, Michael J. Greco, and Christopher J. Petrone. Exploring trends in abundance of young-of-the-year and age-1 Atlantic croaker, black drum, spot, and weakfish in relation to salinity, temperature, and large-scale climatic signals in a Mid-Atlantic estuary. Transactions of the American Fisheries Society, 151(2):150–171, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Miranda:2021:HAT

[MLF21] L. E. Miranda, Kurt M. Lakin, and Nicky M. Faucheux. Habitat associations of three black bass species in a reservoir system. Transactions of the American Fisheries Society, 150(4): 538–547, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moynes:2020:CBW

[MLZ^{+}20] Emily N. Moynes, Michael J. Lawrence, Aaron J. Zolderdo, Tanya S. Prystay, Alice E. I. Abrams, Peter Holder, and Steven J. Cooke. Comparative behavior of wild bluegill captured inside and outside of a long-standing aquatic protected area. Transactions of the American Fisheries Society, 149(2): 174–183, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Molina-Moctezuma:2021:MSD

[MMPZ21] Alejandro Molina-Moctezuma, Erin Peterson, and Joseph D. Zydlewski. Movement, survival, and delays of Atlantic salmon smolts in the Piscataquis River, Maine, USA. Transactions of the American Fisheries Society, 150(3):345–360, May 2021. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:2023:CTG

[MOE^{+}23] Benjamin E. Martin, Brian O’Malley, Randy L. Eshenroder, Yu-Chun Kao, Chris M. Olds, Timothy P. O’Brien, and Chris L. Davis. Comparison of traditional and geometric morphometrics using Lake Huron ciscoes of the Coregonus artedi complex. Transactions of the American Fisheries Society, 152(3): 296–309, May 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mullen:2020:AES

[MSH^{+}20] Eric J. Mullen, Alexandra N. Schoen, McKenzie D. Hauger, Lee Murray, and W. Gary Anderson. Angler experience and seasonal effects on the response of the lake sturgeon to catchand-release angling. Transactions of the American Fisheries Society, 149(6):709–720, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marjomaki:2023:MYW

[MVK23] Timo J. Marjomäki, Pentti Valkeajärvi, and Juha Karjalainen. Males die young, which may cause an Allee effect during a population collapse of the vendace Coregonus albula. Transactions of the American Fisheries Society, 152(6):711–725, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:2023:LCI

[MWS^{+}23] Benjamin E. Meyer, Mark S. Wipfli, Erik R. Schoen, Daniel J. Rinella, and Jeffrey A. Falke. Landscape characteristics influence projected growth rates of stream-resident juvenile salmon in the face of climate change in the Kenai River watershed, south-central Alaska. Transactions of the American Fisheries Society, 152(2):169–186, March 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nepal:2020:DDM

[NF20] Vaskar Nepal and Mary C. Fabrizio. Density-dependence mediates the effects of temperature on growth of juvenile blue catfish in nonnative habitats. Transactions of the American Fisheries Society, 149(1):108–120, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Niles:2021:RTH

[NH21] Jonathan M. Niles and Kyle J. Hartman. Riparian timber harvest intensity affects diets of Appalachian brook trout. Transactions of the American Fisheries Society, 150(4):490–500, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nielsen:2021:WMF

[Nie21] Larry A. Nielsen. Want to make fisheries more diverse and inclusive? Tell new stories! Transactions of the American Fisheries Society, 150(1):3–7, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:2022:RPS

[NJM22] Lindsey Noel Nelson, Christian M. Jones, and Jan R. McDowell. Rangewide population structure of the clearnose skate. Transactions of the American Fisheries Society, 151(3):356–372, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Nathan:2020:EGS

[NKL^{+}20] Lucas R. Nathan, Yoichiro Kanno, Benjamin H. Letcher, Amy B. Welsh, Andrew R. Whiteley, and Jason C. Vokoun. Evaluation of genetic structuring within GIS-derived brook trout management units. Transactions of the American Fisheries Society, 149(6):681–694, November 2020. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:2021:EAL

[NMG^{+}21] Thomas Reid Nelson, Cyril J. Michel, Meagan P. Gary, Brendan M. Lehman, Nicholas J. Demetras, Jeremy J. Hammen, and Michael J. Horn. Effects of artificial lighting at night on predator density and salmonid predation. Transactions of the American Fisheries Society, 150(2):147–159, March 2021. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olmi-Graham:2022:ESL

[OGD22] Helen D. Olmi-Graham and M. Zachary Darnell. Evidence for sperm limitation in the Louisiana blue crab spawning stock. Transactions of the American Fisheries Society, 151(4):399– 407, July 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olmi-Graham:2023:STV

[OGD23] Helen D. Olmi-Graham and M. Zachary Darnell. Spatial and temporal variation in exploitation rates of the Louisiana blue crab spawning stock. Transactions of the American Fisheries Society, 152(3):287–295, May 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:2022:SAL

[OGT22] Zachary Olsen, Ethan Getz, and Perry Trial. Seasonal abundance and length patterns of brown and white shrimp in Texas estuaries and nearshore waters: Implications for population and fishery dynamics. Transactions of the American Fisheries Society, 151(2):215–230, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Owens:2022:QFS

[OK22] Hope W. Owens and Ernest R. Keeley. Quantifying the foodscape for stream-dwelling cutthroat trout reveals spatial and temporal ranges of resource exploitation and energy intake. Transactions of the American Fisheries Society, 151(4):441– 452, July 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olsen:2022:GDD

[OWB^{+}22] Jeffrey B. Olsen, Patrick Walsh, Anna-Marie Benson, Craig J. Schwanke, and John K. Wenburg. Genetic divergence and diversity reflect a predominant freshwater resident life history in rainbow trout from Southwestern Alaska. Transactions of the American Fisheries Society, 151(2):202–214, March 2022. CO-DEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Pennock:2023:CST

[PARM23] Casey A. Pennock, Brandon Albrecht, Ronald J. Rogers, and Mark C. McKinstry. Can spatial and temporal differences in fish assemblage structure inform conservation of an endangered sucker in a large western reservoir? Transactions of the American Fisheries Society, 152(4):475–489, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearsons:2023:DSM

[PGT23] Todd N. Pearsons, Peter J. Graf, and Timothy N. Taylor. Distribution and straying of minijack Chinook salmon released from a captive broodstock hatchery program. Transactions of the American Fisheries Society, 152(4):397–414, July 2023. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Poole:2023:RIM

[PKZW23] Alex S. Poole, Todd M. Koel, Alexander V. Zale, and Molly A. H. Webb. Rotenone induces mortality of invasive lake trout and rainbow trout embryos. Transactions of the American Fisheries Society, 152(1):3–14, January 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:2021:SFL

[PLRR21] Jessica L. Pierce, Matthew V. Lauretta, Ryan J. Rezek, and Jennifer S. Rehage. Survival of Florida largemouth bass in a coastal refuge habitat across years of varying drying severity.

Transactions of the American Fisheries Society, 150(4):435– 451, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearsons:2023:SCH

[PM23] Todd N. Pearsons and Mark D. Miller. Stray compositions of hatchery-origin Chinook salmon and steelhead in natural spawning populations of the upper Columbia watershed. Transactions of the American Fisheries Society, 152(5):515–529, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pennock:2020:RSM

[PMG20] Casey A. Pennock, Mark C. McKinstry, and Keith B. Gido. Razorback sucker movement strategies across a river–reservoir habitat complex. Transactions of the American Fisheries Society, 149(5):620–634, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearsons:2020:SRN

[PO20] Todd N. Pearsons and Rolland R. O’Connor. Stray rates of natural-origin Chinook salmon and steelhead in the Upper Columbia River watershed. Transactions of the American Fisheries Society, 149(2):147–158, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pope:2021:JCS

[PPH^{+}21] Adam C. Pope, Russell W. Perry, Brett N. Harvey, Dalton J. Hance, and Hal C. Hansel. Juvenile Chinook salmon survival, travel time, and floodplain use relative to riverine channels in the Sacramento–San Joaquin River Delta. Transactions of the American Fisheries Society, 150(1):38–55, January 2021. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Porreca:2020:DPS

[PRB^{+}20] Anthony P. Porreca, Neil P. Rude, Kyle J. Broadway, David H. Wahl, and Joseph J. Parkos. Differential performance of sympatric Pomoxis species is explained by habitat-linked morphology. Transactions of the American Fisheries Society, 149(5): 565–575, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulson:2020:DDE

[PS20] Brandon L. Paulson and Craig A. Stockwell. Density-dependent effects of invasive red swamp crayfish Procambarus clarkii on experimental populations of the Amargosa pupfish. Transactions of the American Fisheries Society, 149(1):84–92, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pease:2020:SVT

[PSBGDRH20] Allison A. Pease, Miriam Soria-Barreto, Alfonso A. González-Díaz, and Rocío Rodiles-Hernández. Seasonal variation in trophic diversity and relative importance of basal resources supporting tropical river fish assemblages in Chiapas, Mexico. Transactions of the American Fisheries Society, 149(6): 753–769, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Puncher:2021:TGF

[PWM^{+}21] Gregory Neils Puncher, Yanjun Wang, Ryan Martin, Gregory DeCelles, Steven X. Cadrin, Douglas Zemeckis, Sherrylynn Rowe, Nathalie M. Leblanc, Genevieve J. Parent, and Scott A. Pavey. Transborder gene flow between Canada and the USA and fine-scale population structure of Atlantic cod in the broader Gulf of Maine region. Transactions of the American Fisheries Society, 150(5):560–577, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pritt:2020:VRI

[PZTC20] Jeremy J. Pritt, Richard D. Zweifel, Stephen M. Tyszko, and Joseph D. Conroy. Variation in reproductive investment among Ohio reservoir largemouth bass populations. Transactions of the American Fisheries Society, 149(5):552–564, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Quinn:2021:RAS

[QPS^{+}21] Thomas P. Quinn, George R. Pess, Ben J. G. Sutherland, Samuel J. Brenkman, Ruth E. Withler, Kelsey Flynn, and Terry D. Beacham. Resumption of anadromy or straying? Origins of sockeye salmon in the Elwha River. Transactions of the American Fisheries Society, 150(4):452–464, July 2021. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Railsback:2022:WWD

[Rai22] Steven F. Railsback. What we don’t know about the effects of temperature on salmonid growth. Transactions of the American Fisheries Society, 151(1):3–12, January 2022. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:2021:LAL

[RAZ^{+}21] Connor H. Reid, Alice E. I. Abrams, Aaron J. Zolderdo, Jonathan D. Midwood, E. Don Stevens, Thomas W. Moon, and Steven J. Cooke. A local analgesic, lidocaine, did not affect short-term welfare during electroanesthesia of a teleost fish. Transactions of the American Fisheries Society, 150(4): 477–489, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ressel:2020:DLH

[RBS20] Kirsten N. Ressel, Jenefer L. Bell, and Trent M. Sutton. Distribution and life history of spawning capelin in subarctic Alaska. Transactions of the American Fisheries Society, 149(1):43–56, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roni:2023:CME

[RCC^{+}23] Philip Roni, Meghan J. Camp, Kristin Connelly, Kai Ross, and Hans Berge. A comparison of methods for estimating juvenile salmon habitat capacity to assist with restoration planning and evaluation. Transactions of the American Fisheries Society, 152 (2):201–216, March 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reynolds:2021:EBC

[Rey21] James B. Reynolds. Estimating body conductivity and immobilization threshold in fish. Transactions of the American Fisheries Society, 150(5):593–604, September 2021. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:2022:EPS

[RFR^{+}22a] Connor H. Reid, Matthew D. Faust, Graham D. Raby, Travis O. Brenden, Steven J. Cooke, and Christopher S. Vandergoot. Erratum: Postrelease survival and migration behavior of adult walleye following intracoelomic transmitter implantation using two methods of electro-immobilization. Transactions of the American Fisheries Society, 151(6):743–744, November 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:2022:PSM

[RFR^{+}22b] Connor H. Reid, Matthew D. Faust, Graham D. Raby, Travis O. Brenden, Steven J. Cooke, and Christopher S. Vandergoot. Postrelease survival and migration behavior of adult walleye following intracoelomic transmitter implantation using two methods of electro-immobilization. Transactions of the American Fisheries Society, 151(1):100–111, January 2022. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reid:2022:RA

[RFR^{+}22c] Connor H. Reid, Matthew D. Faust, Graham D. Raby, Travis O. Brenden, Steven J. Cooke, and Christopher S. Vandergoot. Reviewer acknowledgments. Transactions of the American Fisheries Society, 151(6):745–746, November 2022. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Railsback:2023:CTR

[RH23] Steven F. Railsback and Bret C. Harvey. Can thermal refuges save salmonids? Simulation of cold-pool benefits to salmonid populations. Transactions of the American Fisheries Society, 152(4):383–396, July 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Railsback:2021:IDL

[RHA21] Steven F. Railsback, Bret C. Harvey, and Daniel Ayllón. Importance of the daily light cycle in population–habitat relations: a simulation study. Transactions of the American Fisheries Society, 150(1):130–143, January 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryan:2023:EEB

[RKAF23] Athena Ryan, John F. Kocik, Ernest J. Atkinson, and Nathan B. Furey. The effects of environmental and biological factors on the length of Atlantic salmon age-1+ parr in three Maine drainages. Transactions of the American Fisheries Society, 152(3):327–345, May 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radford:2021:CFH

[RLMCC21] Dakota S. Radford, Alec R. Lackmann, Cassi J. Moody-Carpenter, and Robert E. Colombo. Comparison of four hard structures including otoliths for estimating age in blue suckers. Transactions of the American Fisheries Society, 150(4): 514–527, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rubenson:2020:TRJ

[RLO20] Erika S. Rubenson, David J. Lawrence, and Julian D. Olden. Threats to rearing juvenile Chinook salmon from nonnative smallmouth bass inferred from stable isotope and fatty acid biomarkers. Transactions of the American Fisheries Society, 149(3):350–363, May 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ransbury:2022:MCS

[RMS^{+}22] Shane R. Ransbury, Scott Maclean, Corey J. Schwanke, Tracy Hansen, and James W. Savereide. Modeling Chinook salmon run timing in the Gulkana River, Alaska using environmental indices. Transactions of the American Fisheries Society, 151(5): 559–570, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rosinski:2020:NPA

[RVY20] Caroline L. Rosinski, Mark R. Vinson, and Daniel L. Yule. Niche partitioning among native ciscoes and nonnative rainbow smelt in Lake Superior. Transactions of the American Fisheries Society, 149(2):184–203, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ritter:2020:GUR

[RZGL20] Thomas David Ritter, Alexander V. Zale, Grant Grisak, and Michael J. Lance. Groundwater upwelling regulates thermal hydrodynamics and salmonid movements during high-temperature events at a montane tributary confluence. Transactions of the American Fisheries Society, 149(5):600–619, September 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schmitt:2021:EFC

[SHO21] Joseph D. Schmitt, Corbin D. Hilling, and Donald J. Orth. Estimates of food consumption rates for invasive blue catfish. Transactions of the American Fisheries Society, 150(4):465–476, July 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Storch:2021:CAA

[SHS^{+}21] Adam J. Storch, Steven L. Haeseker, Gabriel Scheer, Jerry A. McCann, Brandon R. Chockley, Timothy Copeland, and Robert B. Lessard. Comment: Associations among fish length, dam passage history, and survival to adulthood in two at-risk species of Pacific salmon. Transactions of the American Fisheries Society, 150(2):189–195, March 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic). See response [FBW^{+}21].

Smith:2021:NPK

[SM21] Geoffrey H. Smith and Debra J. Murie. Nonnative pike killifish do not exert predatory pressure on juvenile common snook. Transactions of the American Fisheries Society, 150(6):694– 706, November 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Su:2022:PGA

[SMC^{+}22] Yingxin Su, Peter B. Moyle, Matthew A. Campbell, Amanda J. Finger, Sean M. O’Rourke, Jason Baumsteiger, and Michael R. Miller. Population genomic analysis of the Speckled Dace species complex identifies three distinct lineages in California. Transactions of the American Fisheries Society, 151(6): 695–710, November 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stobart:2022:FSB

[SMM^{+}22] Ben Stobart, Stephen Mayfield, Richard McGarvey, Lisa Carlin, and Julian Morison. “Fishing Smart”: Bioeconomic optimization of fishing strategies in an Australian mollusk fishery. Transactions of the American Fisheries Society, 151(4):408–421, July 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sylvia:2020:ILB

[SMW20] Andrea Sylvia, Brandon Maahs, and Michael J. Weber. Influence of largemouth bass behaviors, angler behaviors, and environmental conditions on fishing tournament capture success. Transactions of the American Fisheries Society, 149(1):93–107, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:2023:BBI

[SN23] William E. Smith and Matthew L. Nobriga. A bioenergeticsbased index of habitat suitability: Spatial dynamics of foraging constraints and food limitation for a rare estuarine fish. Transactions of the American Fisheries Society, 152(5):650– 671, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Snyder:2022:ANF

[SOK^{+}22] Claire E. Snyder, Devon C. Oliver, Brent C. Knights, Stephen M. Pescitelli, and Gregory W. Whitledge. Assessment of native fish passage through Brandon Road Lock and Dam, Des Plaines River, Illinois, using fin ray microchemistry. Transactions of the American Fisheries Society, 151(2):172– 184, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Schaeffer:2020:MSE

[SPVM20] Erin M. Schaeffer, Jeramy J. Pinkerton, Paul A. Venturelli, and Loren M. Miller. Muskellunge spatial ecology in the St. Louis River Estuary and Southwestern Lake Superior. Transactions of the American Fisheries Society, 149(6):651–663, November 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Small:2020:GEP

[SSM20] Hamish J. Small, Jingwei Song, and Jan R. McDowell. Genetic evaluation of population structure in Tautog. Transactions of the American Fisheries Society, 149(4):414–426, July 2020. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scarnecchia:2022:PSR

[SSS^{+}22] Dennis L. Scarnecchia, Jason D. Schooley, Aaron Slominski, K. Michael Backes, Brandon Brown, Brent D. Gordon, and Youngtaik Lim. Patterns and scaling of reproductive output in paddlefish Polyodon spathula and its conservation implications. Transactions of the American Fisheries Society, 151(3): 347–355, May 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Sweka:2022:ISE

[SW22] John A. Sweka and Tyler Wagner. Influence of seasonal extreme flows on brook trout recruitment. Transactions of the American Fisheries Society, 151(2):231–244, March 2022. CODEN ???? ISSN 0363-2415 (print), 1548-8446 (electronic).

Taylor:2020:FSS

[Tay20] Eric B. Taylor. Fishes of the Salish Sea. Transactions of the American Fisheries Society, 149(2):241, March 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thalmann:2020:TAW

[TDB20] Hillary L. Thalmann, Elizabeth A. Daly, and Richard D. Brodeur. Two anomalously warm years in the Northern California Current: Impacts on early marine steelhead diet composition, morphology, and potential survival. Transactions of the American Fisheries Society, 149(4):369–382, July 2020. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teletchea:2020:RBT

[Tel20] Fabrice Teletchea. Reproductive biology of teleost fishes. Transactions of the American Fisheries Society, 149(1):131, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tucker:2022:IHO

[THH^{+}22] Stefan Tucker, Christopher J. Houghton, Brandon Harris, Robert Elliott, Michael Donofrio, and Patrick S. Forsythe. Identification of high-occupancy areas and movement dynamics of age-0 lake sturgeon in the Lower Fox River, Wisconsin. Transactions of the American Fisheries Society, 151(6): 682–694, November 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Triano:2022:AEP

[TKM^{+}22] Ben Triano, Kevin M. Kappenman, Thomas E. McMahon, Matt Blank, Kurt C. Heim, Albert E. Parker, Alexander V. Zale, Nolan Platt, and Katey Plymesser. Attraction, entrance, and passage efficiency of Arctic grayling, trout, and suckers at Denil fishways in the Big Hole River Basin, Montana. Transactions of the American Fisheries Society, 151(4):453–473, July 2022. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taylor:2022:IHA

[TM22] Christopher M. Taylor and Kevin B. Mayes. Impact of hydrologic alteration on Brazos River pelagophilic minnows. Transactions of the American Fisheries Society, 151(4):474–486, July 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tattam:2020:STI

[TR20] Ian A. Tattam and James R. Ruzycki. Smolt transportation influences straying of wild and hatchery Snake River steelhead into the John Day River. Transactions of the American Fisheries Society, 149(3):284–297, May 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tillotson:2023:EBC

[TWP23] Nathan A. Tillotson, Michael J. Weber, and Clay L. Pierce. Effects of bigheaded carp on larval freshwater drum diets. Transactions of the American Fisheries Society, 152(5):530– 549, September 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Voss:2022:IPT

[VACS^{+}22] Nicholas S. Voss, Robert Al-Chokhachy, Adam J. Sepulveda, Christine E. Verhille, Michael P. Ruggles, and Alexander V. Zale. The importance of phenology and thermal exposure to early life history success of nonnative smallmouth bass in the Yellowstone River. Transactions of the American Fisheries Society, 151(5):527–542, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Voss:2023:RSA

[VBQ23] Nicholas S. Voss, Brett J. Bowersox, and Michael C. Quist. Reach-scale associations between introduced brook trout and juvenile and stream-resident bull trout in Idaho. Transactions of the American Fisheries Society, 152(6):835–848, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vacura:2023:WMA

[VCV23] Kirsten Vacura, Edward Camp, and Paul Venturelli. What makes anglers happy: a sentiment analysis of walleye angler forums in the United States. Transactions of the American Fisheries Society, 152(6):749–759, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vinson:2023:LSK

[VHA^{+}23] Mark R. Vinson, Matthew E. Herbert, Amanda S. Ackiss, Jamie A. Dobosenski, Lori M. Evrard, Owen T. Gorman, Joshua F. Lyons, Sydney B. Phillips, and Dan L. Yule. Lake Superior Kiyi reproductive biology. Transactions of the American Fisheries Society, 152(1):75–93, January 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanderBloemen:2023:IIB

[VMS^{+}23] Spencer VanderBloemen, Leandro E. Miranda, Greg G. Sass, Michael Colvin, and Nicky Faucheux. Influence of invasive bigheaded carps on abundance of gizzard shad in the Tennessee River. Transactions of the American Fisheries Society, 152(6): 809–818, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walsh:2020:PFI

[Wal20] Stephen J. Walsh. Proceedings of the First International Snakehead Symposium. Transactions of the American Fisheries Society, 149(3):364–365, May 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winsor:2021:PMS

[WBAB21] Stephen Winsor, Steve Blumenshine, Paul Adelizi, and Matt Bigelow. Precocious maturation in spring-run Chinook salmon is affected by incubation temperature, feeding regime, and parentage. Transactions of the American Fisheries Society, 150 (5):578–592, September 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weitkamp:2022:LFL

[WBD^{+}22] Laurie A. Weitkamp, Brian R. Beckman, Donald M. Van Doornik, Angelica Munguia, Mary Hunsicker, and Meredith Journey. Life in the fast lane: Feeding and growth of juvenile steelhead and Chinook salmon in main-stem habitats of the Columbia River Estuary. Transactions of the American Fisheries Society, 151(5):587–610, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wojtusik:2023:ROA

[WBKK23] Kristopher J. Wojtusik, David L. Berlinsky, Linas W. Kenter, and Adrienne I. Kovach. River-of-origin assignment of migratory striped bass, with implications for mixed-stock analysis. Transactions of the American Fisheries Society, 152(1):15–34, January 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Welsh:2020:TSS

[WCSJ20] Amy B. Welsh, Douglas M. Carlson, Scott L. Schlueter, and James R. Jackson. Tracking stocking success in a long-lived species through genetics and demographics: Evidence of natural reproduction in lake sturgeon after twenty-two years. Transactions of the American Fisheries Society, 149(1):121–130, January 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

White:2023:BRI

[Whi23] J. Wilson White. Book review: Introduction to Quantitative Ecology. Timothy E. Essington. Oxford University Press, Oxford, UK. 2021. 304 pages. $97.00 (hardcover), $48.99 (paperback), $17.27 (ebook). Transactions of the American Fisheries Society, 152(6):849–850, November 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wippelhauser:2021:RDF

[Wip21] Gail Wippelhauser. Recovery of diadromous fishes: a Kennebec River case study. Transactions of the American Fisheries Society, 150(3):277–290, May 2021. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallerius:2020:HAI

[WJCA20] Magnus Lovén Wallerius, Jörgen I. Johnsson, Steven J. Cooke, and Robert Arlinghaus. Hook avoidance induced by private and social learning in common carp. Transactions of the American Fisheries Society, 149(4):498–511, July 2020. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolters-Rinker:2022:ACS

[WRRG22] Pilar N. Wolters-Rinker, David L. Rogowski, and Alice C. Gibb. Allometric changes and shape comparisons of two southwestern sucker species and their hybrids using a geometric morphometric approach. Transactions of the American Fisheries Society, 151(5):630–640, September 2022. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:2022:SPA

[WSS^{+}22] Sterling Watson, Abraham Schneider, Leon Santen, Katherine A. Deters, Robert Mueller, Brett Pflugrath, John Stephenson, and Z. Daniel Deng. Safe passage of American eels through a novel hydropower turbine. Transactions of the American Fisheries Society, 151(6):711–724, November 2022. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Williams:2023:VME

[WZS^{+}23] Benjamin Williams, Shalima Zalsha, S. Lynne Stokes, Ryan P. A. McShane, and John Foster. Validating matches of electronically reported fishing trips to investigate matching error. Transactions of the American Fisheries Society, 152(1):114– 126, January 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yard:2023:DPP

[YYK^{+}23] Michael D. Yard, Charles B. Yackulic, Josh Korman, Michael J. Dodrill, and Bridget R. Deemer. Declines in prey production during the collapse of a tailwater rainbow trout population are associated with changing reservoir conditions. Transactions of the American Fisheries Society, 152(1):35–50, January 2023. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).